### v1

# Language Modeling for the Future of Finance: A Quantitative Survey into Metrics, Tasks, and Data Opportunities

Abstract
Recent advances in language modeling have led to growing interest in applying Natural Language Processing (NLP) techniques to financial problems, enabling new approaches to analysis and decision-making. To systematically examine this trend, we review 374 NLP research papers published between 2017 and 2024 across 38 conferences and workshops, with a focused analysis of 221 papers that directly address finance-related tasks. We evaluate these papers across 11 qualitative and quantitative dimensions, identifying key trends such as the increasing use of general-purpose language models, steady progress in sentiment analysis and information extraction, and emerging efforts around explainability and privacy-preserving methods. We also discuss the use of evaluation metrics, highlighting the importance of domain-specific ones to complement standard machine learning metrics. Our findings emphasize the need for more accessible, adaptive datasets and highlight the significance of incorporating financial crisis periods to strengthen model robustness under real-world conditions. This survey provides a structured overview of NLP research applied to finance and offers practical insights for researchers and practitioners working at this intersection.
1 Introduction
Language modeling has become a key tool in natural language processing (NLP) for analyzing unstructured text, including reports, news articles, and social media posts. These methods are increasingly applied to finance-related tasks such as sentiment analysis (Balakrishnan et al., 2022), information extraction (Huang et al., 2023), summarization (Khanna et al., 2022), stock prediction (Jain & Agrawal, 2024), and volatility forecasting (Niu et al., 2023). Given the increasing number of NLP papers addressing financial problems, there is a need to systematically examine how the field is contributing to finance-related applications. While prior surveys have examined this intersection (Table 1), many adopt a qualitative approach (Chen et al., 2022b; Gao et al., 2021b; Xiao et al., 2024a), focusing on broad NLP techniques (Jagdale & Deshmukh, 2025; Man et al., 2019; Liu, 2024), deep learning (Ozbayoglu et al., 2020) and Large Language Models (LLMs) (Nie et al., 2024; Li et al., 2023c), or specific tasks such as sentiment analysis (Mishev et al., 2020). However, these studies often lack systematic evaluation.
In this study, we focus specifically on NLP research applied to finance. Our scope includes papers published in NLP venues where methods are developed, benchmarked, or adapted for financial problems. We analyze 374 papers published from 2017 to 2024 across 38 NLP conferences and workshops. After further filtering (Section 2), we retain 221 papers that directly address finance-related tasks. These are evaluated across 11 qualitative and quantitative dimensions, including tasks, methodologies, datasets, evaluation metrics, and accessibility.
Our analysis reveals several patterns in the current research landscape. Commonly addressed tasks include sentiment analysis, information extraction, and question answering, while areas such as explainability and privacy-preserving techniques are explored less frequently (Section 3). Standard machine learning metrics are widely used for evaluation, although these may not always reflect domain-specific priorities (Section 7.1). Most studies focus on stable market conditions, with less attention to more volatile periods (Section 7.2) and relying on outdated sources or overlooking issues such as survivorship bias (Section 7.3) Finally, we observe a trend toward the use of general-purpose language models over custom architecture (Section 4). While this reflects the rapid integration of recent advances in NLP, further exploration of domain-specific strategies may offer additional value for finance-related tasks.
Paper Review Years Papers Reviewed Features Area surveyed Systematic Collection Domain trends Quantitative Analysis Temporal Analysis
(Gao et al., 2021a) 1959 (A)-2020 87 5 General Overview × × × ×
(Liu, 2024) 2022-2024 49 11 General Overview × × × ×
(Millo et al., 2024) 2018-2023 30 1 Methodologies × × ×
(Chen et al., 2020b) 2016-2019 62 3 Financial Technology × × × ×
(Chen et al., 2022a) 2018-2022 38 2 Financial Technology × × ×
(Xing et al., 2017) 1998 (A)-2016 127 4 Financial Forecasting × × ×
(Zhao et al., 2024a) 2004 (A)-2024 146 1 Large Language Models × × ×
(Li et al., 2024) 2020-2023 68 1 Large Language Models × × × ×
(Dong et al., 2024) 2023-2024 206 2 Large Language Models × ×
(Nie et al., 2024) 2019-2024 318 1 Large Language Models × × ×
(Lee et al., 2025) 2018-2023 51 3 Large Language Models × × × ×
(Man et al., 2019) 2004-2019 89 1 Machine Learning × × × ×
(Ozbayoglu et al., 2020) 1998 (A)-2020 151 6 Deep Learning × ×
(Mishev et al., 2020) 2003 (A)-2020 89 1 Sentiment Analysis × ×
Language Modeling for the Future of Finance 2017-2024 374 11 Tasks, Methodologies, Data, Metrics, Code, Authorship, Funding
Table 1: Comparison of previous Natural Language Processing surveys in finance based on their focus areas, number of papers reviewed, and analytical features. Due to the absence of systematic collection methods in most prior research work, entries marked with (A) in the "Review Years" column indicate that the reported range is approximate.
2 Paper Extraction Process
Figure 1: Overview of our paper selection process and analysis dimensions. We collected papers from a broad range of NLP venues using abstract-level keyword filtering, yielding 374 candidates. After removing mismatches and shared task papers, we retained 221 papers, categorized into four groups based on their connection to financial tasks. The analysis focuses on the selected dimensions as shown in (Table 5, Appendix B).
To study how NLP has been applied to finance, we selected papers from 38 NLP venues, including CoLM, ACL, NAACL, EMNLP, LREC, COLING, and workshops like FinNLP and the Workshop on Economics and Natural Language Processing. The full list of venues is provided in Table 9 (Appendix E). As illustrated in Figure 1, we began by filtering papers that mentioned “finance” or “financial” in their abstracts. Out of all such papers published from 1975 to the present, more than 94% were published from 2017 onward. This growth coincides with the emergence of transformer-based models (Vaswani et al., 2017), which greatly expanded the scope of NLP applications. Accordingly, we use 2017 as a threshold year for our analysis.
Category Description
Financial Forecast (Category I) Targets predicting financial events, including stock movements, volatility, bankruptcy, and currency exchange rates.
Financial Resources (Category II) Covers tasks addressing finance-specific issues beyond prediction, such as constructing financial datasets, detecting fraud in finance-related documents, and extracting financial events.
Financial Applications (Category III) Focuses on general ML/NLP tasks like sentiment analysis and information/relation extraction for financial datasets.
Finance Related (Category IV) Covers tasks not applied to financial data or targeting financial problems, but potentially useful in finance, such privacy-preserving and explainable AI.
Table 2: Summary of task categories based on the specifics of their financial focus, ranging from direct prediction of financial outcomes to general NLP tasks with potential financial value.
This initial filtering returned 374 papers. Among them, 88 were tied to shared tasks (e.g., SemEval-2017 (Kar et al., 2017), FinCausal-2022 (Mondal et al., 2022)), and 65 used the term “financial” in a different contexts – such as referencing financial resources (Sekeres et al., 2024; Ding & Riloff, 2018) – without any actual financial application. After removing these, we retained 221 papers for our analysis.
Unlike previous surveys, which often emphasize qualitative observations, our study combines both qualitative and quantitative methods to provide a broader view of the field. We categorized the selected papers by their primary focus, resulting in four task-based categories (Figure 2). This classification helps us examine trends in how NLP techniques are being used to tackle finance-relevant problems and supports more detailed comparisons across types of tasks.
3 Task Distribution in NLP for Financial Applications
Figure 2: Distribution of primary tasks across categories. Each cell shows the task name and paper count (e.g., "Sentiment Analysis (3)"), with color gradients indicating frequency – darker shades represent more papers. "Miscellaneous" groups tasks that appear only once within Categories II and III.
To understand how NLP has been used in financial contexts, we categorized papers into four groups (shown in Table 2) based on their primary tasks (Figure 2). This classification helps map how language modeling techniques are applied and how directly each paper engages with finance, with Category I representing the most direct financial use cases. Detailed descriptions of each category are provided in Appendix D.
Financial Forecast (Category I) papers cover predictive tasks such as stock price and volatility forecasting. While these areas are well-studied, others like economic forecasting (Arno et al., 2023), risk assessment (Zhou et al., 2020), and cryptocurrency prediction (Seroyizhko et al., 2022) receive less focus. These areas present opportunities for expanding the reach of prediction models and improving their robustness and comparability (Section 7).
Financial Resources (Category II) papers often focus on dataset construction (Section 4, Figure 3). These datasets – ranging from annotated earnings calls to news and speeches – support tasks like financial event extraction (Huang et al., 2024; Ju et al., 2023), fraud detection (Erben & Waldis, 2024; Wang et al., 2019), and annotation (Aguda et al., 2024; Khatuya et al., 2024). Although these tasks contribute to enhancing financial data processing, other tasks remain relatively underrepresented.
In the Financial Applications (Category III) group, sentiment analysis (Rodriguez Inserte et al., 2023), information extraction (Lior et al., 2024), and question answering (Kosireddy et al., 2024) are the dominant tasks. These methods are widely used to analyze earnings calls, reports, and market commentary, where investor sentiment and factual extraction are key inputs for decision-making. Recent efforts have aimed to improve QA models for handling financial text (Theuma & Shareghi, 2024; Liu et al., 2024; Mavi et al., 2023).
Finally, among Finance-Related (Category IV) papers, the most studied areas are dataset construction and numerical reasoning. While these datasets are not finance-specific, they have potential applications in finance. For instance, fake news detection datasets (Vargas et al., 2021) can help reduce misinformation in markets. Numerical reasoning (Akhtar et al., 2023), important for understanding financial statements and assessing risk, is receiving more attention. However, other areas – such as explainable AI (XAI) (Klein & Walther, 2024) and privacy-preserving methods (Abbe et al., 2012) – remain rarely explored, despite their relevance to secure and interpretable decision-making (Basu et al., 2021).
4 Language Model Adoption & Datasets
Figure 3: Timeline of PLM/LLM adoption in NLP research applied to finance, alongside key datasets by task type: C (custom models), QA (question answering), IE (information extraction), OL (other language resources), and FD (financial documents). For each model, the top date indicates release; the bottom, first usage in the surveyed papers.
Acceleration of Model Adoption
Figure 3 (Language Models) shows how quickly language models are adopted for finance-related NLP tasks. While early models like BERT (Devlin et al., 2019) took 13 months to appear in these applications, recent models – GPT-3.5, GPT-4 (Achiam et al., 2023), and Mistral (Jiang et al., 2023) – were adopted within 6–12 months. LLaMA-2 (Touvron et al., 2023)) saw integration in just 5 months, reflecting the community’s increasing speed in applying general NLP advances to financial tasks.
Dominance of General-Purpose Models
As shown in Figure 3 (Language Models), general-purpose models – LLMs like GPT-4 and LLaMA-2, and Pretrained Language Models (PLMs) like BERT and RoBERTa (Liu et al., 2019)) – are used far more frequently than domain-adapted ones. BERT appears in over 70 papers, while finance-tuned models such as FinBERT (Araci, 2019; Yang et al., 2020b; Liu et al., 2020; Huang et al., 2022), FinMA (Xie et al., 2024a), and FinGPT (Yang et al., 2023) are less common. Notably, earlier FinBERT versions are more widely used than later ones, suggesting that improvements in general models often outweigh the benefits of domain-specific adaptation.
Decline of Custom Architectures
Custom models were common before 2022, especially for tasks like table-based QA (e.g., TagOp for TAT-QA (Zhu et al., 2021), HyBrider for HybridQA (Chen et al., 2020c)). Some models, like MDRM (Qin & Yang, 2019), HTML (Yang et al., 2020a), and HAN (Hu et al., 2021), are still used as baselines, particularly in volatility prediction (Niu et al., 2023; Mathur et al., 2022). However, the rise of general LLMs has led to a sharp decline in custom architectures. As a result of the accelerated development and deployment of these general purpose models, researchers increasingly rely on adapting pretrained models instead of building domain-specific ones from scratch.
Language Coverage in Datasets
The OL and FD sections of Figure 3 highlights the scarcity of non-English datasets for forecasting tasks such as return or volatility prediction. One exception is CMIN (Luo et al., 2023), which includes English and Chinese financial data. Most datasets use English-only sources like EDGAR filings (Maheshwari et al., 2022) and financial news (Xie et al., 2024b), which may limit model performance in non-English markets. Creating multilingual forecasting datasets would support broader applicability and reduce language bias, as also noted by Jørgensen et al. (2023).
Since 2017, research applying NLP to finance has shifted from custom models toward general-purpose pretrained architectures. New models are adopted faster than before, and general PLMs have largely replaced finance-specific ones. This preference also influences dataset use: most forecasting datasets are English-only, limiting generalization to global markets.
5 Methods Used in NLP for Finance
Figure 4: Trends in methodology adoption over time, showing stable use of traditional techniques and rapid growth in PLMs and LLMs. Less frequent methods are omitted for clarity.
To understand the methodological landscape, we grouped the approaches used in the surveyed papers into nine categories (Appendix C, Table 6). While the total number of papers has grown, methods like statistical NLP, conventional machine learning, and embeddings have remained mostly flat. In contrast, there has been a sharp increase in the use of deep learning, PLMs and LLMs, reflecting the field’s shift toward pretrained architectures.
Shift Toward PLMs and LLMs
As seen in Figure 4, the first three categories – statistical NLP, embeddings, and conventional ML – have persisted but have seen reduced usage as PLMs and LLMs take lead. The rise began with models like BERT in 2019, which significantly improved context handling in financial texts (Li et al., 2017; Chen et al., 2019; Ein-Dor et al., 2019). PLMs are now widely used for tasks such as summarization and entity recognition. Since 2023, LLMs like ChatGPT-3.5 have seen rapid adoption (Tang et al., 2023; Li et al., 2023a; Ni et al., 2023), especially for text generation (Guo et al., 2023b) and financial QA (Uluoglakci & Temizel, 2024). This transition has also brought more complex deep learning architectures used for multi-step reasoning and domain-specific challenges (Ma et al., 2024; Zhao et al., 2024b; Su et al., 2024).
Underexplored Techniques
While pretrained models dominate, several methods remain underused despite their promise. Statistical modeling, graph-based approaches, and knowledge graphs could improve model robustness and interpretability, particularly for applications like causality detection (Tabari et al., 2018) or market dynamics (Menzio et al., 2024). Emerging techniques like GraphRAG (Edge et al., 2024) highlight the potential of such methods, yet they appear in only a fraction of papers. Their limited presence suggests opportunities for expanding methodological diversity in future research.
6 Code & Data Accessibility
Figure 5: Temporal trends in code and dataset availability, highlighting the shift toward open-source practices and the growing accessibility of NLP resources for finance.
As the adoption of language models accelerates, the issue of open accessibility in NLP research, especially in finance, becomes more relevant. As shown in Figure 5, open-source practices have become more common in recent years. Before 2021, closed-source code dominated, but the trend has shifted toward sharing code and datasets. This move enhances transparency, reproducibility, and collaborative progress in NLP research applied to finance (Whited, 2023). Despite this, some papers still include inactive links or empty repositories, limiting reproducibility. At the same time, dataset availability has steadily increased since 2017, pointing to growing awareness around open data. Maintaining functional, up-to-date repositories remains essential to support meaningful benchmarking and model development.
7 Potential in Financial Forecasting
Figure 6: Distribution of evaluation metrics used in Category I papers. Most rely on ML-based metrics, while only a few financial metrics appear repeatedly.
Category I papers reveal several areas where forecasting models could be improved. Most studies use general ML metrics, leaving out finance-specific measures that better capture practical utility. Crisis periods are rarely considered, limiting insights into how models behave under stress. Finally, there is a strong U.S. and English-language bias in data, with limited adoption of global financial datasets. Expanding these areas of exploration could strengthen model robustness and support broader real-world applicability.
7.1 Potential in Financial Metrics
As shown in Figure 6, Category I papers mostly rely on ML metrics like accuracy, F1, or MSE (Sawhney et al., 2020b; Wu, 2020; Yangjia et al., 2022). While useful, these do not fully reflect financial performance. Only three financial metrics appear more than once – Sharpe Ratio, Maximum Drawdown, and Cumulative Return (defined in Appendix G).
Finance-specific metrics offer risk-adjusted insights that ML measures overlook. Sharpe Ratio, for example, evaluates return relative to risk (Zou et al., 2022). Incorporating such metrics would improve comparability across models and align research outcomes with practical needs (Zhang et al., 2024; Zou et al., 2022; Sawhney et al., 2021b).
7.2 Crisis Years for Stress Testing
Figure 7: Distribution of data years used in Category I papers, highlighting publication trends, recession periods, and key financial data milestones.
Figure 7 shows that data usage begins to rise after 1993, when electronic filings became publicly available, with another increase in 2005 following the HTML filing mandate (U.S. Securities and Exchange Commission, ). However, usage dips around 2009 and 2020 – years associated with the global financial crisis and COVID-19 pandemic. These periods are essential for evaluating model robustness, yet they remain underused, likely due to data availability challenges prior to 2015 (Bloomberg L.P., 2015).
Despite these constraints, crisis periods offer valuable testing grounds. Most models are trained on stable conditions, which may not reflect their behavior during volatility. Including data from crises enables stress testing (Investopedia Staff, 2025b) and helps build forecasting models that remain reliable under instability.
7.3 Data Sources
Commonly Used and Overlooked Data Sources
Most forecasting models use common financial sources: stock prices, SEC filings, financial news, earnings calls (e.g., MDRM (Qin & Yang, 2019), StockNet (Xu & Cohen, 2018)). But many valuable datasets are either rarely used (e.g., Federal Reserve reports (Menzio et al., 2024)) or completely absent from NLP research. Public resources like FRED (Federal Reserve Bank of St. Louis, 2024) and Fama-French (Fama & French, 2024) offer rich macroeconomic indicators that remain unexplored. Incorporating these datasets could help capture macroeconomic trends and market dynamics, improving model reliability.
Data Accessibility and Maintenance
Access limitations are also a concern. Legacy datasets like MDRM are outdated, and APIs such as Yahoo Finance’s have been discontinued. Others, like CRSP, require paid access. This limits reproducibility and restricts access to reliable, up-to-date data. There’s a clear need for open, regularly updated datasets applicable for financial forecasting.
8 Discussion & Conclusion
This paper presents a systematic, quantitative analysis of NLP research applied to finance. From an initial pool of 374 papers across major NLP venues (Appendix E), we curated a final set of 221 papers after filtering. We categorized them by primary task (Section 3) and assessed each across 11 qualitative and quantitative dimensions (Table 5). Table 3 summarizes key trends and identifies opportunities where language modeling research can further contribute to finance-related problems.
Criteria Trends and Observations Potential Opportunities and Recommendations
NLP Tasks in Finance Sentiment analysis, information extraction, and question answering are the most frequently addressed tasks. Forecasting is mostly centered around stock and volatility prediction. Underexplored areas such as explainability, privacy-preserving methods, and tasks like bankruptcy or cryptocurrency prediction present valuable directions for future work (Section 3).
PLM/LLM Adoption General-purpose PLMs and LLMs have largely replaced finance-specific and custom architectures. At the same time, the use of statistical NLP and conventional ML continues to decline (Section 4). Revisiting finance-specific models and exploring alternative methods such as graph-based learning or hybrid statistical models could offer improvements for financial tasks (Section 5).
Evaluation Metrics Most studies rely on ML metrics such as accuracy and MSE, which do not reflect financial performance. Incorporating financial metrics like Sharpe Ratio or Maximum Drawdown would improve the practical relevance and comparability of predictive models (Section 7.1).
Crisis Periods Few studies include data from financial crises such as 2008–2009 or 2020–2021, focusing instead on stable periods. Including data from volatile periods can support stress-testing and help build models more resilient to real-world fluctuations (Section 7.2).
Data Selection and Bias Most studies use English-language datasets and U.S.-based financial sources. While there is growing diversity in the types of data used (e.g., news, filings, social media), much of it is static, with limited updates or adaptation to changing financial contexts. Developing multilingual and globally representative datasets, along with maintaining and updating existing ones, would support better generalization and long-term applicability of models across financial domains (Section 7.3).
Open Accessibility Code and dataset sharing has become more common, especially after 2021, enhancing transparency and reproducibility. Open access should be a standard in NLP-for-finance research when legally feasible. Maintaining functional and up-to-date repositories is key for long-term reproducibility (Section 5).
Table 3: Summary of trends and research directions in NLP research applied to finance, highlighting areas of focus, methodological shifts, and key opportunities for future exploration.
Implications for NLP & LM Researchers
NLP research applied to finance has largely centered on a few core tasks – sentiment analysis, information extraction, and question answering – while underexploring others like explainability and privacy-preserving methods (Section 3). The rapid adoption of PLMs and LLMs has made it easier to apply language models, but has also led to reduced methodological variety (Section 4). Graph-based models, statistical techniques, and other alternatives remain underused despite their potential (Section 5). While open-source practices have become more common, researchers should be more careful about maintaining functional and up-to-date repositories to ensure long-term reproducibility and meaningful benchmarking (Section 5). In addition, adopting financial-domain metrics can make evaluations more meaningful and aligned with practical goals (Section 7.1). Including crisis periods in datasets (Section 7.2) and improving data selection practices (Section 7.3) can help ensure robustness and consistency. More diverse datasets, especially beyond English and U.S. markets, are also needed to broaden applicability (Section 7.3).
Considerations for Financial Practitioners
While NLP methods are increasingly used in financial contexts, several challenges remain. Many models are trained on static or incomplete datasets, leading to biases that may affect predictions (Section 7.3). Evaluations often omit real-world considerations such as market volatility and financial risk (Sections 7.2, 7.1). Incorporating financial benchmarks and testing models under stress conditions could improve their practical value. Continued progress in this direction can better align research outcomes with decision-making needs in financial settings.
Ethics Statement
This study does not assess the merit or quality of individual papers. We do not suggest that any category or method is inherently better than others. Our goal is to map the volume and distribution of research applying NLP methods to finance, and to offer a foundation for further study, without making judgments about the value of specific approaches.
Acknowledgments
We did not receive any specific funding for this work. We thank Arnav Hiray and Michael Galarnyk for their comments.
Appendix A Authorship and Funding
Based on the analysis we conducted, we also made a binary classification of the following categories:
  • Authors affiliated with academia
  • Authors affiliated with Industry
  • Funding from industry
  • Funding from government sources
  • Funding from academic sources
Figure 8: This figure displays the correlation between authors from industrial and academic backgrounds within various types of funding such as academic, industrial and governmental.
As seen in the figure 8, we see that there is no major correlation between funding and authorship. Thus it does not imply that if an author with either industrial or academic is present in the list of authors, there will necessarily be funding from the industry/academia/government.
Category Count
Papers with Academic and Industrial authors 78
Papers with only Academic authors 118
Papers with only Industrial authors 25
Papers with Industrial Funding 49
Papers with Governmental Funding 61
Papers with Academic Funding 35
Table 4: Summary of the count of authors and funding within each category.
Appendix B Various Dimensions surveyed
Exploration dimension Description
Primary task The main task addressed in the paper, such as volatility prediction (Wang et al., 2024).
Sub-Tasks Additional tasks that support the primary objective, like using sentiment analysis to enhance stock price predictions (Jain & Agrawal, 2024).
Methodology Techniques applied, ranging from traditional machine learning to deep learning and large language models. See Section 5.
Code and data availability Whether the paper provides open-source code or data, and whether those artifacts are accessible (e.g., active links).
Contribution What type of contribution the paper has such as a dataset, framework, model, evaluation, and so on.
Comparison/baselines The baselines the paper compares their work to.
Data source The types of data used, including financial reports, social media, news articles, or other financial documents.
Metrics Evaluation criteria, including standard ML metrics (accuracy, precision) and finance-specific metrics (Sharpe ratio).
Data years The time periods covered by the data, such as crisis years or more stable financial periods.
Authorship Indicates what affiliation the authors of the paper have (academic, industrial).
Funding source The origin of the research funding (academic, industrial, or governmental) if any.
Table 5: This table outlines the dimensions examined in our analysis of NLP research focused on the financial domain. Dimensions that led to remarkable findings are discussed in the main paper, while an analysis of authorship and funding is provided in Appendix A.
Appendix C Common Methods used
Methodology Description
Statistical NLP Methods like TF-IDF, Bag-of-Words, and n-grams focused on extracting statistical patterns from financial text.
Embeddings Techniques such as Word2Vec, GloVe, and custom embeddings designed to map financial terms into vector spaces, improving downstream tasks.
Conventional ML Algorithms like Logistic Regression, Support Vector Machines (SVM), and Decision Trees, often used for classification and risk prediction.
Deep Learning (DL) Custom neural network architectures tailored for specific financial tasks such as stock price forecasting.
PLMs Encoder-only models (e.g., BERT) and Encoder-Decoder models (e.g., T5) are used for tasks like document summarization and information extraction.
LLMs Models like GPT-3.5 and LLaMA-2 focus on text generation and understanding complex financial language.
Statistical Modeling Involves correlation analysis, Granger causality, etc., to understand relationships in financial data.
Graphs Uses graph structures to model interactions in financial systems.
Knowledge Graphs Integrating structured knowledge within financial tasks to enhance model performance and reliability.
Table 6: This table categorizes the key methodologies into nine groups.“PLMs” stands for Pretrained Language Models, while “LLMs” stands for Large Language Models.
Appendix D Task Description
The following Table 7 and Table 8 describe the task description with examples.
Primary Task Examples & Insights
Finance-Related (Categories IV)
Explainable AI (XAI) refers to methods that make machine learning outputs more interpretable and less of a "black box." These papers show how small input changes can drastically affect model explanations (Sinha et al., 2021) and highlight the instability of tools like LIME in sensitive applications (Burger et al., 2023). In finance, where decisions must be traceable and justifiable, clear and reliable explanation methods are essential for building trust in tasks like fraud detection and risk assessment.
Privacy-preserving methods protect sensitive data while still allowing model training and analysis. Examples include homomorphic encryption for text similarity (Kim et al., 2022), federated learning for distributed training without raw data exchange (Zhao et al., 2024d), and domain adaptation that shares only model parameters (Xiao et al., 2024b). These methods are especially important in finance, where institutions must work with private data while staying compliant and avoiding breaches.
Numerical reasoning involves working with numbers and structured data to solve problems or develop algorithms. This includes solving math word problems using declarative knowledge (Roy & Roth, 2018), learning numeracy through number embeddings (Duan et al., 2021), and pretraining models for verifying tabular claims (Akhtar et al., 2023). In finance, good numerical reasoning supports tasks like market prediction and risk evaluation by improving how models interpret and reason about numbers.
Financial Applications (Category III)
Sentiment analysis helps assess market mood and forecast financial movements using sources like news, social media, and reports. Studies analyze opinions in reports and news (Rodriguez Inserte et al., 2023), investor sentiment on social media (Guo et al., 2023a), and tone in financial texts (Choe et al., 2023). As markets become more influenced by public opinion, sentiment analysis plays a growing role in trading and analysis.
Information extraction (IE) and relation extraction focus on identifying entities and linking them to events or other entities. Examples include extracting financial events from documents (Zheng et al., 2019), signals in reports (Huang et al., 2023), and patterns in social media (Conforti et al., 2022). These tools help connect companies, individuals, and financial instruments to real-world developments (Liou et al., 2021), supporting real-time analysis and prediction.
Question Answering (QA) enables systems to find specific information in financial documents, reports, and databases. QA models support quick access to relevant facts (Mavi et al., 2023), and recent work focuses on calibrating large language models (LLMs) for finance-specific tasks (Zhao et al., 2024c; Theuma & Shareghi, 2024; Addlesee, 2024). This area emphasizes not only building QA tools but also improving their accuracy for use in high-stakes financial contexts.
Table 7: Overview of key primary tasks in Categories IV (Finance-Related) above the mid line and Category III (Financial Applications) below. Each entry summarizes the task’s role and relevance in NLP research applied to finance, with representative examples and practical insights.
Primary Task Examples & Insights
Financial Resources (Category II)
Dataset and Resource Construction involves building structured, labeled financial datasets for model development and evaluation. These papers focus on creating large-scale annotated resources (Chen et al., 2020a), including datasets like Tab-CQA (Liu et al., 2023) and ConvFinQA (Chen et al., 2022c) designed for reasoning over tables and multi-step numerical queries. Other works assemble corpora from Earnings calls (Pardawala et al., 2025), financial reports (Shah et al., 2022), company filings (Zmandar et al., 2022), news (Tang et al., 2023), and government documents (Shah et al., 2023), enabling downstream tasks like entity linking, numerical reasoning, or forecasting.
Fraud detection leverages domain-specific data and models to identify deceptive financial behavior across platforms. Erben & Waldis (2024) identifies financial scams on Instagram using a fine-tuned BERT model deployed via browser extension and REST API. Another approach detects identity fraud through interactive dialogue, employing knowledge graphs and reinforcement learning to expose inconsistencies in claimed personal data (Wang et al., 2019). These systems highlight how NLP architectures can protect users from financial manipulation.
Financial Forecast (Category I)
Stock Price and Volatility Prediction aims to forecast stock movements or market instability using text data. These models use historical data (Sawhney et al., 2021a), news (Ahbali et al., 2022), and event-driven signals (Sawhney et al., 2020a; Wu, 2020) to predict stock trends. Volatility prediction further explores market sensitivity by analyzing unstructured text like press releases or financial articles (Qin & Yang, 2019), helping traders anticipate fluctuations.
Risk Prediction involves identifying the likelihood and impact of adverse financial events, such as defaults or market disruptions. These models analyze unstructured data like earnings call transcripts (Sang & Bao, 2022), regulatory filings, and legal documents to detect early signals of financial risk (Li et al., 2023b). Applications include credit risk estimation, fraud detection, and systemic risk monitoring (Zhang et al., 2024).
Table 8: Descriptions of major primary tasks in Category II (Financial Resources) above the line and Category I (Financial Forecast) below. Each entry summarizes the task’s focus and contribution to NLP research applied to finance, with representative examples.
Appendix E Conference Proceedings
Category Conferences and Workshops
Financial Forecast (Category I) ACL, Australasian Language Technology Association Workshop, CCL, EACL, EcoNLP Workshop, EMNLP, EMNLP (Industry), Financial Technology and Natural Language Processing (FinNLP) Workshop, ICCL, LREC-COLING, NAACL, NAACL (Industry), SRW
Financial Resources (Category II) ACL, ACL (Industry), CoNLL, EACL, EcoNLP Workshop, EMNLP, EMNLP-IJCNLP, e-Commerce and NLP Workshop, Financial Narrative Processing and MultiLing Financial Summarisation Workshop, Financial Technology and Natural Language Processing (FinNLP) Workshop, GWC, ICON, ICCL, LREC, LREC-COLING, NAACL, NAACL (Industry), NLP4PI Workshop, Pattern-based Approaches to NLP in the Age of Deep Learning Workshop, SIGHUM Workshop
Financial Applications (Category III) ACL, ATALA, Bridging Human–Computer Interaction and Natural Language Processing Workshop, CCL, CoLM, Computational Approaches to Subjectivity, Sentiment and Social Media Analysis Workshop, DeeLIO Workshop, EACL, EcoNLP Workshop, EMNLP, EMNLP (Industry), EMNLP-IJCNLP, Financial Narrative Processing and MultiLing Financial Summarisation Workshop, Financial Technology and Natural Language Processing (FinNLP) Workshop, ICON, ICCL, INLG, LREC, LREC-COLING, NAACL, NAACL (Industry), Natural Legal Language Processing Workshop, News Media Content Analysis and Automated Report Generation Workshop, Pattern-based Approaches to NLP in the Age of Deep Learning Workshop, Safety4ConvAI Workshop, Structured Prediction for Natural Language Processing Workshop, TextGraphs
Finance Related (Category IV) ACL, ACL-IJCNLP, BlackboxNLP Workshop, EMNLP, Financial Technology and Natural Language Processing (FinNLP) Workshop, LREC-COLING, SRW, TrustNLP Workshop
Table 9: Conference and Workshop Categorization
Appendix F Papers per year
Year 2024 2023 2022 2021 2020 2019 2018 2017 2016 2015 2014 2012 2010 2008 2006 2002 2001 1976 1975
Number of Papers 127 53 90 32 57 15 14 21 6 1 5 1 3 3 1 1 1 1 1
Table 10: Number of papers satisfying our filtration criterion per year.
Appendix G Glossary
Metric Definition
Sharpe Ratio Measures how much return an investment gives for each unit of risk, compared to a risk-free asset (Investopedia Staff, 2025a).
Maximum Drawdown The biggest drop from a peak to a low point in a portfolio’s value before it recovers (Investopedia Staff, 2023b).
Cumulative Return The total profit or loss from an investment over time (Investopedia Staff, 2023a).
Table 11: Key Definitions of financial terms used in our paper.

### v2

@@
-Language Modeling for the Future of Finance: A Quantitative Survey into Metrics, Tasks, and Data Opportunities
+Language Modeling for the Future of Finance: A Survey into Metrics, Tasks, and Data Opportunities
 Abstract
@@
 To systematically examine this trend, we review 374 NLP research papers published between 2017 and 2024 across 38 conferences and workshops, with a focused analysis of 221 papers that directly address finance-related tasks.
-We evaluate these papers across 11 qualitative and quantitative dimensions, identifying key trends such as the increasing use of general-purpose language models, steady progress in sentiment analysis and information extraction, and emerging efforts around explainability and privacy-preserving methods.
-We also discuss the use of evaluation metrics, highlighting the importance of domain-specific ones to complement standard machine learning metrics.
-Our findings emphasize the need for more accessible, adaptive datasets and highlight the significance of incorporating financial crisis periods to strengthen model robustness under real-world conditions.
-This survey provides a structured overview of NLP research applied to finance and offers practical insights for researchers and practitioners working at this intersection.
+We evaluate these papers across 11 quantitative and qualitative dimensions, and our study identifies the following opportunities: (i) expanding the scope of forecasting tasks; (ii) enriching evaluation with financial metrics; (iii) leveraging multilingual and crisis-period datasets; and (iv) balancing PLMs with efficient or interpretable alternatives.
+We identify actionable directions for research and practice, supported by dataset and tool recommendations, with implications for both the academia and industry communities.
 1 Introduction
 Language modeling has become a key tool in natural language processing (NLP) for analyzing unstructured text, including reports, news articles, and social media posts.
-These methods are increasingly applied to finance-related tasks such as sentiment analysis (Balakrishnan et al., 2022), information extraction (Huang et al., 2023), summarization (Khanna et al., 2022), stock prediction (Jain & Agrawal, 2024), and volatility forecasting (Niu et al., 2023).
+These methods are increasingly applied to finance-related tasks such as sentiment analysis (Balakrishnan et al., 2022), information extraction (Huang et al., 2023), summarization (Khanna et al., 2022), stock prediction (Jain and Agrawal, 2024), and volatility forecasting (Niu et al., 2023).
 Given the increasing number of NLP papers addressing financial problems, there is a need to systematically examine how the field is contributing to finance-related applications.
-While prior surveys have examined this intersection (Table 1), many adopt a qualitative approach (Chen et al., 2022b; Gao et al., 2021b; Xiao et al., 2024a), focusing on broad NLP techniques (Jagdale & Deshmukh, 2025; Man et al., 2019; Liu, 2024), deep learning (Ozbayoglu et al., 2020) and Large Language Models (LLMs) (Nie et al., 2024; Li et al., 2023c), or specific tasks such as sentiment analysis (Mishev et al., 2020).
-However, these studies often lack systematic evaluation.
+While prior surveys have examined this intersection (Table 1), many lack a structured approach (Chen et al., 2022a; Gao et al., 2021b; Xiao et al., 2024), focusing on broad NLP techniques (Jagdale and Deshmukh, 2025; Man et al., 2019; Liu, 2024), deep learning (Ozbayoglu et al., 2020) and Large Language Models (LLMs) (Nie et al., 2024; Li et al., 2023), or specific tasks such as sentiment analysis (Mishev et al., 2020).
 In this study, we focus specifically on NLP research applied to finance.
@@
 After further filtering (Section 2), we retain 221 papers that directly address finance-related tasks.
-These are evaluated across 11 qualitative and quantitative dimensions, including tasks, methodologies, datasets, evaluation metrics, and accessibility.
-Our analysis reveals several patterns in the current research landscape.
-Commonly addressed tasks include sentiment analysis, information extraction, and question answering, while areas such as explainability and privacy-preserving techniques are explored less frequently (Section 3).
-Standard machine learning metrics are widely used for evaluation, although these may not always reflect domain-specific priorities (Section 7.1).
-Most studies focus on stable market conditions, with less attention to more volatile periods (Section 7.2) and relying on outdated sources or overlooking issues such as survivorship bias (Section 7.3) Finally, we observe a trend toward the use of general-purpose language models over custom architecture (Section 4).
-While this reflects the rapid integration of recent advances in NLP, further exploration of domain-specific strategies may offer additional value for finance-related tasks.
+These are evaluated across 11 quantitative and qualitative dimensions, including tasks, methodologies, datasets, evaluation metrics, and accessibility.
+Table 1: Comparison of previous Natural Language Processing surveys in finance based on their focus areas, number of papers reviewed, and analytical features.
+Due to the absence of systematic collection methods in most prior research work, entries marked with (A) in the "Review Years" column indicate the approximate range.
 Paper Review Years Papers Reviewed Features Area surveyed Systematic Collection Domain trends Quantitative Analysis Temporal Analysis
@@
 (Millo et al., 2024) 2018-2023 30 1 Methodologies × × ×
-(Chen et al., 2020b) 2016-2019 62 3 Financial Technology × × × ×
-(Chen et al., 2022a) 2018-2022 38 2 Financial Technology × × ×
+(Chen et al., 2020a) 2016-2019 62 3 Financial Technology × × × ×
+(Chen et al., 2022b) 2018-2022 38 2 Financial Technology × × ×
 (Xing et al., 2017) 1998 (A)-2016 127 4 Financial Forecasting × × ×
-(Zhao et al., 2024a) 2004 (A)-2024 146 1 Large Language Models × × ×
+(Zhao et al., 2024) 2004 (A)-2024 146 1 Large Language Models × × ×
 (Li et al., 2024) 2020-2023 68 1 Large Language Models × × × ×
@@
 Language Modeling for the Future of Finance 2017-2024 374 11 Tasks, Methodologies, Data, Metrics, Code, Authorship, Funding
-Table 1: Comparison of previous Natural Language Processing surveys in finance based on their focus areas, number of papers reviewed, and analytical features.
-Due to the absence of systematic collection methods in most prior research work, entries marked with (A) in the "Review Years" column indicate that the reported range is approximate.
+Our analysis reveals not only valuable insights but also actionable directions for both research and practice with dataset and tool recommendations.
+Forecasting tasks, while well-established, leave room for exploration, especially in areas such as risk and macroeconomic prediction (Section 3).
+Financial evaluation metrics are gaining traction and could further improve the practical applicability of models (Section 4.1).
+The growing availability of temporally diverse (Sections 4.2, 4.3), multilingual and multimodal (Sections 4.4) datasets enables more robust, globally applicable models.
+Finally, while PLMs and LLMs dominate, lightweight and interpretable methods remain viable for practical deployment (Section 5).
 2 Paper Extraction Process
-Figure 1: Overview of our paper selection process and analysis dimensions.
-We collected papers from a broad range of NLP venues using abstract-level keyword filtering, yielding 374 candidates.
-After removing mismatches and shared task papers, we retained 221 papers, categorized into four groups based on their connection to financial tasks.
-The analysis focuses on the selected dimensions as shown in (Table 5, Appendix B).
-To study how NLP has been applied to finance, we selected papers from 38 NLP venues, including CoLM, ACL, NAACL, EMNLP, LREC, COLING, and workshops like FinNLP and the Workshop on Economics and Natural Language Processing.
-The full list of venues is provided in Table 9 (Appendix E).
+To study how NLP has been applied to finance, we selected papers from 38 NLP venues, including ACL, NAACL, EMNLP, LREC, CoLM, COLING, and workshops like FinNLP and the Workshop on Economics and Natural Language Processing.
 As illustrated in Figure 1, we began by filtering papers that mentioned “finance” or “financial” in their abstracts.
@@
 Accordingly, we use 2017 as a threshold year for our analysis.
+The keyword filtering approach enabled us to cast a wide net, allowing for a high volume of papers.
+We were then able to filter these further using manual methods.
+This initial filtering returned 374 papers.
+Among them, 88 were tied to shared tasks (e.g., SemEval-2017 (Kar et al., 2017), FinCausal-2022 (Mondal et al., 2022)), and 65 used the term “financial” in a different contexts (such as referencing financial resources (Sekeres et al., 2024; Ding and Riloff, 2018)) without any actual financial application.
+After removing these, we retained 221 papers for our analysis.
+Unlike previous surveys, which often emphasize qualitative observations, our study combines quantitative and qualitative methods for a broader view of the field.
+We categorized the selected papers by their primary focus into four task-based categories (Figure 2).
+This classification helps us examine trends in how NLP techniques are used to tackle finance-relevant problems and supports more detailed comparisons across types of tasks.
+3 Task Distribution in NLP for Financial Applications
+To understand how NLP has been used in financial contexts, we categorized papers into four distinct groups (shown in Table 2) based on their primary tasks (Figure 2).
+This classification provides a mapping from language modeling techniques to their applications and the level of engagement with finance, with Category I papers representing direct financial use cases.
+Table 2: Summary of task categories based on the specifics of their financial focus, ranging from direct prediction of financial outcomes to general NLP tasks with potential financial value.
 Category Description
@@
 Finance Related (Category IV) Covers tasks not applied to financial data or targeting financial problems, but potentially useful in finance, such privacy-preserving and explainable AI.
-Table 2: Summary of task categories based on the specifics of their financial focus, ranging from direct prediction of financial outcomes to general NLP tasks with potential financial value.
-This initial filtering returned 374 papers.
-Among them, 88 were tied to shared tasks (e.g., SemEval-2017 (Kar et al., 2017), FinCausal-2022 (Mondal et al., 2022)), and 65 used the term “financial” in a different contexts – such as referencing financial resources (Sekeres et al., 2024; Ding & Riloff, 2018) – without any actual financial application.
-After removing these, we retained 221 papers for our analysis.
-Unlike previous surveys, which often emphasize qualitative observations, our study combines both qualitative and quantitative methods to provide a broader view of the field.
-We categorized the selected papers by their primary focus, resulting in four task-based categories (Figure 2).
-This classification helps us examine trends in how NLP techniques are being used to tackle finance-relevant problems and supports more detailed comparisons across types of tasks.
-3 Task Distribution in NLP for Financial Applications
-Figure 2: Distribution of primary tasks across categories.
-Each cell shows the task name and paper count (e.g., "Sentiment Analysis (3)"), with color gradients indicating frequency – darker shades represent more papers.
-"Miscellaneous" groups tasks that appear only once within Categories II and III.
-To understand how NLP has been used in financial contexts, we categorized papers into four groups (shown in Table 2) based on their primary tasks (Figure 2).
-This classification helps map how language modeling techniques are applied and how directly each paper engages with finance, with Category I representing the most direct financial use cases.
-Detailed descriptions of each category are provided in Appendix D.
+Figure 1: Overview of our paper selection process and analysis dimensions.
+We collected papers from a broad range of NLP venues using abstract-level keyword filtering, yielding 374 candidates.
+After removing mismatches and shared task papers, we retained 221 papers, categorized into four groups by their connection to financial tasks.
 Financial Forecast (Category I) papers cover predictive tasks such as stock price and volatility forecasting.
-While these areas are well-studied, others like economic forecasting (Arno et al., 2023), risk assessment (Zhou et al., 2020), and cryptocurrency prediction (Seroyizhko et al., 2022) receive less focus.
-These areas present opportunities for expanding the reach of prediction models and improving their robustness and comparability (Section 7).
-Financial Resources (Category II) papers often focus on dataset construction (Section 4, Figure 3).
-These datasets – ranging from annotated earnings calls to news and speeches – support tasks like financial event extraction (Huang et al., 2024; Ju et al., 2023), fraud detection (Erben & Waldis, 2024; Wang et al., 2019), and annotation (Aguda et al., 2024; Khatuya et al., 2024).
-Although these tasks contribute to enhancing financial data processing, other tasks remain relatively underrepresented.
+While these tasks are well-studied, areas like economic forecasting (Arno et al., 2023), risk assessment (Zhou et al., 2020), and cryptocurrency prediction (Seroyizhko et al., 2022) receive less focus.
+These areas offer opportunities to expand the reach of predictive models, improving robustness and comparability (Section 4).
+Financial Resources (Category II) papers often focus on dataset construction (Figure 6).
+Ranging from annotated earnings calls to news and speeches, These datasets support tasks like financial event extraction (Huang et al., 2024; Ju et al., 2023), fraud detection (Erben and Waldis, 2024; Wang et al., 2019), and annotation (Aguda et al., 2024; Khatuya et al., 2024).
+While these tasks enhance financial data processing, other tasks remain relatively underrepresented.
 In the Financial Applications (Category III) group, sentiment analysis (Rodriguez Inserte et al., 2023), information extraction (Lior et al., 2024), and question answering (Kosireddy et al., 2024) are the dominant tasks.
 These methods are widely used to analyze earnings calls, reports, and market commentary, where investor sentiment and factual extraction are key inputs for decision-making.
-Recent efforts have aimed to improve QA models for handling financial text (Theuma & Shareghi, 2024; Liu et al., 2024; Mavi et al., 2023).
+Recent efforts have aimed to improve QA models for handling financial text (Theuma and Shareghi, 2024; Liu et al., 2024; Mavi et al., 2023).
 Finally, among Finance-Related (Category IV) papers, the most studied areas are dataset construction and numerical reasoning.
@@
 Numerical reasoning (Akhtar et al., 2023), important for understanding financial statements and assessing risk, is receiving more attention.
-However, other areas – such as explainable AI (XAI) (Klein & Walther, 2024) and privacy-preserving methods (Abbe et al., 2012) – remain rarely explored, despite their relevance to secure and interpretable decision-making (Basu et al., 2021).
-4 Language Model Adoption & Datasets
-Figure 3: Timeline of PLM/LLM adoption in NLP research applied to finance, alongside key datasets by task type: C (custom models), QA (question answering), IE (information extraction), OL (other language resources), and FD (financial documents).
+However, other areas – such as explainable AI (XAI) (Klein and Walther, 2024) and privacy-preserving methods (Abbe et al., 2012) – remain rarely explored, despite their relevance to secure and interpretable decision-making (Basu et al., 2021).
+Figure 2: Distribution of primary tasks across categories.
+Each cell shows the task name and paper count (e.g., "Sentiment Analysis (3)"), with color gradients indicating frequency – darker shades represent more papers.
+"Miscellaneous" groups tasks that appear only once within Categories II and III.
+4 Potential in Financial Forecasting
+Category I papers reveal several areas where forecasting models could be improved.
+Most studies use general ML metrics, leaving out finance-specific measures that better capture practical utility (Section  4.1).
+Crisis periods are rarely considered, limiting insights into how models behave under stress (Section  4.2).
+Finally, there is a strong U.S.
+and English-language bias in data, with limited adoption of global financial datasets (Section  4.3,  4.4).
+Expanding these areas of exploration could strengthen model robustness and support broader real-world applicability.
+4.1 Advancing Evaluation with Finance Specific Metrics
+As shown in Figure 3, Category I papers mostly rely on ML metrics such as accuracy, F1, or MSE (Sawhney et al., 2020; Wu, 2020; Yangjia et al., 2022).
+While useful, these do not fully reflect financial performance.
+Finance-specific metrics offer risk-adjusted insights that ML measures overlook.
+Incorporating such metrics would improve comparability across models and align research outcomes with practical needs (Zhang et al., 2024; Zou et al., 2022; Sawhney et al., 2021).
+Consider two models that predict daily stock returns.
+Model A achieves a lower mean squared error (MSE) than Model B, yet when deployed in a trading strategy, it results in frequent trades, high volatility, and poor performance during market downturns.
+In contrast, Model B, despite its higher MSE, generates a smoother equity curve, experiences lower drawdown, and delivers a higher Sharpe Ratio thus making it more appealing from an application perspective.
+This highlights a critical insight emphasized in financial machine learning literature: traditional statistical metrics like MSE are often inadequate for evaluating trading strategies, as they fail to capture real-world profitability and risk exposure (Bailey et al., 2015; de Prado, 2018).
+Instead, financial metrics such as Sharpe Ratio, Maximum Drawdown, and Cumulative Return are essential because they reflect the model’s risk-adjusted performance and robustness across market regimes (Lo, 2002; Krauss et al., 2017; Takahashi et al., 2009) as well as the applicability of the model to practical real world situations.
+Key Financial Metrics
+There are certain financial metrics that measure models’ performance in a financial setting and provide insight into the their applicability.
+Sharpe Ratio quantifies risk-adjusted return by comparing excess returns to volatility (Investopedia, 2025a).
+Maximum Drawdown (MDD) captures the largest loss from peak to trough, indicating downside risk and model robustness (Hayes, 2024).
+Cumulative Return measures total profit over a period: a direct benchmark for model outputs (Investopedia, 2007).
+Tools for Metric Calculation
+Several Python libraries support quantitative finance and model-driven investment strategies.
+PyPortfolioOpt (Martin, 2021) offers portfolio optimization tools including Sharpe Ratio and risk-return metrics.
+QuantLib (QuantLib Team, 2000) provides analytics for pricing, drawdown, and custom risk modeling.
+Backtrader (Backtrader Team, 2015) enables backtesting of trading strategies with performance evaluation, ideal for simulating NLP-driven trading.
+bt (Morissette, 2025) supports rapid prototyping and comparison of forecasting models with built-in financial metrics.
+4.2 Strengthening Robustness by Incorporating Crisis Periods
+Figure 4 shows that researchers use financial data dating back to 1993, when electronic filings became publicly available, with another rise in 2005 after the HTML filing mandate (U.S.
+SEC, 2025).
+However, most publications use post-2009 (the global financial crisis) and pre-2020 (the COVID-19 pandemic) data, thus missing the crisis periods in evaluation.
+Crisis periods are essential for evaluating model robustness, allowing stress testing (Investopedia, 2025b) and helping build forecasting models that remain reliable under instability.
+Consider the case of Long-Term Capital Management (LTCM).
+Its quantitative models were highly leveraged and calibrated almost exclusively on stable market data, failing to account for extreme tail events.
+When the 1997 Asian and 1998 Russian crises struck, LTCM’s Value-at-Risk (VaR) models severely underestimated exposure, leading to large-scale losses as spreads exploded and liquidity evaporated.
+Within weeks, LTCM lost over 90% of its equity, triggering systemic risk concerns across the global financial system (Department of Land Economy, 2003; Federal Reserve History, 2025; Vanity Fair Staff, 1998).
+As noted in academic and regulatory analyses, this collapse illustrates the dangers of relying solely on backtests from stable periods without incorporating the possibility of rare but catastrophic shocks.
+Most current models are trained in stable conditions that may not reflect their behavior during volatility.
+While there were data availability challenges prior to 2015 (Bloomberg L.P., 2015), the crisis years remain underused.
+In Section 4.3, we highlight valuable data sources, including those with crisis years data.
+4.3 Opportunities in Data Coverage
+Most forecasting models use common financial sources: stock prices, SEC filings, financial news, earnings calls.
+But many valuable datasets are either rarely used (e.g., Federal Reserve reports (Shah et al., 2023; Menzio et al., 2024)) or completely absent from NLP research (e.g., shareholder letters).
+Public resources like FRED (Federal Reserve Bank of St.
+Louis, 2024) and Fama-French (Fama and French, 2024) offer rich but underexplored macro indicators.
+Incorporating these datasets could help capture macroeconomic trends and market dynamics, improving model reliability.
+Below, we highlight sources of valuable financial data, including those already used by the community and the other beneficial ones.
+Figure 3: Distribution of evaluation metrics used in Category I papers. Most rely on ML-based metrics, while only a few financial metrics appear repeatedly.
+Regulatory Filings and Earnings Calls
+U.S.
+public companies file a range of documents such as 10-K, 10-Q, 8-K, credit agreements, proxy statements, S-1, S-3, and others are accessible via the SEC’s EDGAR platform, through free API calls and RSS feeds (U.S.
+SEC, 2024).
+While similar registries exist (e.g., Canada’s SEDAR, Japan’s EDINET, UK’s Companies House), they are rarely used, reinforcing a U.S.-centric and English-language bias (Section 4.4).
+Earnings calls, where firms discuss quarterly performance with investors, are another rich data source.
+Seeking Alpha transcribes  4,500 calls quarterly (Seeking Alpha, 2025); other providers include The Motley Fool (The Motley Fool, 2025), FactSet (FactSet Insight, 2025), Thomson Reuters (Thomson Reuters, 2025), and S&P Global (S&P Global Market Intelligence, 2025), though often behind paywalls.
+Companies also post transcripts/slides on IR websites.
+Public datasets include transcriptions and audio, such as S&P 500 calls from 2017 (Qin and Yang, 2019) and 1,213 companies’ data covering 2015-2018 (Li et al., 2020a).
+Figure 4: Data year distribution in financial forecasting papers, annotated with major financial events and infrastructure milestones.
+Highlights underuse of crisis periods despite their importance for model robustness.
+Financial News and Social Media
+Real-time news from sources like Reuters, Bloomberg, Dow Jones, CNBC, FT, and WSJ is vital for forecasting.
+The Reuters-21578 dataset (UC Irvine, 2023), with 10,369 articles from 1987, remains a benchmark in text analysis.
+Larger archives such as FNSPID (Zhou et al., 2024) link 15.7 million articles (1999-2023) to S&P 500 stocks.
+News is accessible via APIs (e.g., Bloomberg (Bloomberg L.P., 2025a)) and free sources like Yahoo Finance (Yahoo Finance, 2025) and Google News (Google News, 2025).
+On social media, “FinTwit” and platforms like X and StockTwits are used for sentiment analysis.
+Key datasets include Sentiment140 (Kaggle Contributor, 2025; Chaudhary, 2020), StockNet (Xu and Cohen, 2018), and Reddit datasets (Kaggle Contributor, 2021; Wang et al., 2024).
+X’s API offers limited free and full historical access under paid tiers (X (2025), formerly Twitter).
+StockTwits, with Bullish/Bearish sentiment tagging, offers free and commercial API access (StockTwits, 2025) and is often used in forecasting studies.
+Macroeconomic and Monetary Data
+Central bank communications (policy statements, minutes, and press releases) offer critical macroeconomic guidance.
+The U.S.
+Federal Reserve’s FOMC releases detailed statements and transcripts on interest rates, inflation, and economic outlook (Federal Reserve System, 2025), with archives available on the Fed’s website.
+Despite their market relevance, such texts remain underused in NLP-based forecasting.
+Shah et al.
+(2023) addresses this by introducing an annotated dataset of FOMC communications and proposing a sentence-level classification task (hawkish, dovish, neutral) using fine-tuned transformers.
+Economic indicators like GDP, CPI, and unemployment figures – issued by the BEA (U.S.
+BEA, 2025), FRED (Federal Reserve Bank of St.
+Louis, 2024), Eurostat, IMF, OECD, and the World Bank (The World Bank, 2025) – are mostly numerical but often include press commentary, all freely available via official portals.
+Analyst Reports
+These are proprietary equity research reports with buy/sell calls and sector outlooks.
+Data providers such as Thomson Reuters, FactSet, Bloomberg, and S&P I/B/E/S aggregate these reports, typically through subscription services.
+All-in-one platforms like the Bloomberg Terminal (Bloomberg L.P., 2025b) and S&P Capital IQ (S&P Global, 2025) provide access to equity research, earnings estimates, and corporate event data as part of their premium offerings.
+There is no comprehensive free source; however, excerpts or key takeaways occasionally surface in news articles or on platforms like Seeking Alpha.
+4.4 Data Concerns and Biases
+Data Accessibility Concerns
+Despite the presence of a large variety of data sources, there are significant limitations in data accessibility.
+Legacy datasets described in Section 4.3 often become outdated, and APIs for accessing data may be discontinued, as the Yahoo Finance API (EODHD, 2024), or restricted by paywalls.
+Other essential databases, such as CRSP, require paid access and are unavailable to most individual researchers.
+These barriers severely limit reproducibility and restrict access to reliable, up-to-date data.
+As a result, there is a clear need for open, regularly updated financial datasets suitable for forecasting.
+Data Language and Country Bias
+As shown in the previous sections, financial NLP research is heavily skewed toward English-language data, particularly from U.S.
+markets.
+Most forecasting datasets rely on U.S.
+regulatory filings, earnings calls, and financial news, which is often centered on large-cap firms like those in the S&P 500.
+This narrow focus limits the applicability of models to non-English and non-U.S.
+contexts.
+Figure 6 highlights that while some multilingual datasets exist, they are mostly designed for language modeling rather than financial forecasting.
+Very few datasets enable forecasting tasks in other languages, with rare exceptions like the bilingual CMIN corpus (Luo et al., 2023).
+Expanding multilingual, market-diverse datasets is essential for building globally robust forecasting models and mitigating systemic biases in NLP for finance, as also noted by (Jørgensen et al., 2023).
+Multimodal Data Scarcity
+Despite significant progress in multimodal learning, the domain of NLP in Finance still lacks comprehensive datasets that integrate text, audio, and other critical financial modalities.
+One of the few existing resources is MAEC, a dataset aligning transcripts with audio from over 3,400 S&P 500 earnings calls, totaling more than 920 hours of speech, specifically designed for financial risk forecasting (Li et al., 2020b).
+More recently, FinAudio has emerged as the first standardized benchmark focusing on audio-first tasks (ASR on short and long earnings call segments, plus summarization), compiling over 400 hours of financial audio from multiple sources to evaluate Audio-LLM capabilities (Cao et al., 2025).
+While many existing datasets support exploration of acoustic and textual cues in financial forecasting, truly multimodal resources (those integrating visual data such as stock charts, presentation slides alongside structured financial metrics) remain very limited.
+Notable exceptions include Mathur et al.
+(2022b) and Fons et al.
+(2024), which begin to address this gap.
+This modality gap highlights the pressing need for richer, tri- or quad-modal datasets to better capture the multifaceted signals that influence market dynamics.
+5 Broader Insights for NLP in Finance
+5.1 Complementing Foundation Models with Practical Alternatives
+Based on our observations from our corpus of research papers, the methodological landscape in NLP for Finance has shifted decisively toward pretrained language models (PLMs) and large language models (LLMs).
+We group existing approaches into several broad categories, which allows us to track adoption trends and the rise of PLMs, beginning with the use of BERT-based models in finance around 2019 (Araci, 2019a; Yang et al., 2020c), and accelerating rapidly with the emergence of general-purpose LLMs like ChatGPT (Wu et al., 2023; Gao et al., 2023).
+However, NLP tasks in Finance often pose domain-specific constraints such as interpretability, regulatory transparency, data scarcity, and the need for low-latency systems.
+These factors make non-PLM methods (such as rule-based classifiers, feature-engineered models, and domain-specific embeddings) not only viable but often preferable in production environments (Shah et al., 2023; Lopez et al., 2021).
+The rest of this section revisits key methodological alternatives to PLMs, each addressing practical and conceptual needs that remain critical in industrial financial tasks.
+Latency and Scalability
+Classical NLP pipelines are far more resource-efficient than LLMs.
+In time-time contexts like news analytics or alert systems, lightweight models (e.g., logistic regression or small CNNs) deliver insights instantly and can scale to massive corpora without requiring GPU infrastructure (Zhai et al., 2019).
+Systems like RavenPack (Cornez, 2018) and Refinitiv News Analytics (TRNA) (Waeosri, 2020) use rule-based NLP and horizontally scalable architectures to deliver structured news data with sub-second latency, supporting real-time trading and high-throughput analysis.
+Interpretability and Compliance
+Lexicon- and rule-based methods remain essential in regulated settings.
+Financial sentiment lexicons (e.g., Loughran-McDonald) (Loughran and McDonald, 2011) and keyword rules are transparent, auditable, and often outperform neural models in compliance or risk-flagging tasks (Hosseini et al., 2018).
+Lexicon-based surveillance tools, such as those used in communications monitoring (NICE Actimize, 2023) and financial crime compliance (Risk & Compliance, 2021), are widely adopted for their transparency and ease of audit, allowing firms to justify alerts with specific keywords.
+Information Extraction and Knowledge Graphs
+Named-entity and relation extraction systems power knowledge graphs and schema-driven applications.
+These systems offer precision, structure, and explainability, which are ideal for fraud detection, Know Your Customer (KYC), and regulatory alignment, where PLMs often lack controllability (Szarvas et al., 2007; Tarnopolski et al., 2019).
+Financial institutions use rule-based NLP for entity and relation extraction to build and update knowledge graphs in KYC and Anti-Money Laundering (AML) workflows, as seen in solutions from KYCHub (2025) and TigerGraph (2024), enabling automated risk monitoring and explainable link discovery.
+Data Efficiency and Domain Adaptability
+Feature-based pipelines using TF-IDF, static embeddings, and classical classifiers (SVMs, XGBoost) work well with limited data.
+They enable explicit domain knowledge integration (such as tagging financial terms or numerical cues) for credit scoring, bankruptcy prediction, and event detection (Li et al., 2018; Khandani et al., 2021; Alanis et al., 2022).
+Financial firms often use TF-IDF (Simha, 2021) and XGBoost (Kamau, 2024) to build effective models with limited labeled data without requiring costly infrastructure.
+Structure and Summarization
+Unsupervised techniques like LDA and extractive summarization remain valuable for discovering latent topics or summarizing dense financial documents.
+These methods are fast, interpretable, and competitive on factual corpora like earnings calls or filings (Blei et al., 2003; Nenkova and McKeown, 2011).
+Companies like BlackRock, Bank of America (Phoenix Strategy Group, 2025), and American Express (Arora and Radhakrishnan, 2020) apply topic modeling and extractive summarization to earnings calls and filings, using unsupervised or minimally supervised methods for reliable and traceable outputs.
+5.2 Implications of Shift Toward General-Purpose Models
+General-purpose language models have rapidly become the default choice in NLP applied to Finance, with models like RoBERTa (Liu et al., 2019), GPT-4 (Achiam et al., 2023), and LLaMA-2 (Touvron et al., 2023) adopted far more widely and quickly than domain-specific alternatives.
+As Figure 6 shows, even earlier versions of FinBERT (Araci, 2019b; Yang et al., 2020b; Liu et al., 2020; Huang et al., 2022) are used more than newer ones, and recent finance-tuned models like FinGPT (Yang et al., 2023) remain rare.
+This trend reflects how easy it has become to apply strong general models out of the box.
+For practitioners, this means many models are optimized for benchmarks, not for the realities of finance, where interpretability, latency, or regulation often matter more than raw accuracy.
+Figure 5: Trends in code and dataset availability, highlighting the shift toward open-source practices and the growing accessibility of NLP resources for finance.
+Figure 6: Timeline of PLM/LLM adoption in NLP research applied to finance, alongside key datasets by task type: C (custom models), QA (question answering), IE (information extraction), OL (other language), and FD (financial documents).
 For each model, the top date indicates release; the bottom, first usage in the surveyed papers.
-Acceleration of Model Adoption
-Figure 3 (Language Models) shows how quickly language models are adopted for finance-related NLP tasks.
-While early models like BERT (Devlin et al., 2019) took 13 months to appear in these applications, recent models – GPT-3.5, GPT-4 (Achiam et al., 2023), and Mistral (Jiang et al., 2023) – were adopted within 6–12 months.
-LLaMA-2 (Touvron et al., 2023)) saw integration in just 5 months, reflecting the community’s increasing speed in applying general NLP advances to financial tasks.
-Dominance of General-Purpose Models
-As shown in Figure 3 (Language Models), general-purpose models – LLMs like GPT-4 and LLaMA-2, and Pretrained Language Models (PLMs) like BERT and RoBERTa (Liu et al., 2019)) – are used far more frequently than domain-adapted ones.
-BERT appears in over 70 papers, while finance-tuned models such as FinBERT (Araci, 2019; Yang et al., 2020b; Liu et al., 2020; Huang et al., 2022), FinMA (Xie et al., 2024a), and FinGPT (Yang et al., 2023) are less common.
-Notably, earlier FinBERT versions are more widely used than later ones, suggesting that improvements in general models often outweigh the benefits of domain-specific adaptation.
-Decline of Custom Architectures
-Custom models were common before 2022, especially for tasks like table-based QA (e.g., TagOp for TAT-QA (Zhu et al., 2021), HyBrider for HybridQA (Chen et al., 2020c)).
-Some models, like MDRM (Qin & Yang, 2019), HTML (Yang et al., 2020a), and HAN (Hu et al., 2021), are still used as baselines, particularly in volatility prediction (Niu et al., 2023; Mathur et al., 2022).
-However, the rise of general LLMs has led to a sharp decline in custom architectures.
-As a result of the accelerated development and deployment of these general purpose models, researchers increasingly rely on adapting pretrained models instead of building domain-specific ones from scratch.
-Language Coverage in Datasets
-The OL and FD sections of Figure 3 highlights the scarcity of non-English datasets for forecasting tasks such as return or volatility prediction.
-One exception is CMIN (Luo et al., 2023), which includes English and Chinese financial data.
-Most datasets use English-only sources like EDGAR filings (Maheshwari et al., 2022) and financial news (Xie et al., 2024b), which may limit model performance in non-English markets.
-Creating multilingual forecasting datasets would support broader applicability and reduce language bias, as also noted by Jørgensen et al.
-(2023).
-Since 2017, research applying NLP to finance has shifted from custom models toward general-purpose pretrained architectures.
-New models are adopted faster than before, and general PLMs have largely replaced finance-specific ones.
-This preference also influences dataset use: most forecasting datasets are English-only, limiting generalization to global markets.
-5 Methods Used in NLP for Finance
-Figure 4: Trends in methodology adoption over time, showing stable use of traditional techniques and rapid growth in PLMs and LLMs. Less frequent methods are omitted for clarity.
-To understand the methodological landscape, we grouped the approaches used in the surveyed papers into nine categories (Appendix C, Table 6).
-While the total number of papers has grown, methods like statistical NLP, conventional machine learning, and embeddings have remained mostly flat.
-In contrast, there has been a sharp increase in the use of deep learning, PLMs and LLMs, reflecting the field’s shift toward pretrained architectures.
-Shift Toward PLMs and LLMs
-As seen in Figure 4, the first three categories – statistical NLP, embeddings, and conventional ML – have persisted but have seen reduced usage as PLMs and LLMs take lead.
-The rise began with models like BERT in 2019, which significantly improved context handling in financial texts (Li et al., 2017; Chen et al., 2019; Ein-Dor et al., 2019).
-PLMs are now widely used for tasks such as summarization and entity recognition.
-Since 2023, LLMs like ChatGPT-3.5 have seen rapid adoption (Tang et al., 2023; Li et al., 2023a; Ni et al., 2023), especially for text generation (Guo et al., 2023b) and financial QA (Uluoglakci & Temizel, 2024).
-This transition has also brought more complex deep learning architectures used for multi-step reasoning and domain-specific challenges (Ma et al., 2024; Zhao et al., 2024b; Su et al., 2024).
-Underexplored Techniques
-While pretrained models dominate, several methods remain underused despite their promise.
-Statistical modeling, graph-based approaches, and knowledge graphs could improve model robustness and interpretability, particularly for applications like causality detection (Tabari et al., 2018) or market dynamics (Menzio et al., 2024).
-Emerging techniques like GraphRAG (Edge et al., 2024) highlight the potential of such methods, yet they appear in only a fraction of papers.
-Their limited presence suggests opportunities for expanding methodological diversity in future research.
-6 Code & Data Accessibility
-Figure 5: Temporal trends in code and dataset availability, highlighting the shift toward open-source practices and the growing accessibility of NLP resources for finance.
+Table 3: Summary of trends and research directions in NLP research applied to finance, highlighting areas of focus, methodological shifts, and key opportunities for future exploration.
+Criteria Trends and Observations Potential Opportunities and Recommendations
+NLP Tasks in Finance Sentiment analysis, information extraction, and question answering are the most frequently addressed tasks.
+Forecasting is mostly centered around stock and volatility prediction.
+Underexplored areas such as explainability, privacy-preserving methods, and tasks like bankruptcy or cryptocurrency prediction present valuable directions for future work (Section 3).
+Evaluation Metrics Most studies rely on ML metrics such as accuracy and MSE, which do not reflect financial performance.
+Incorporating financial metrics like Sharpe Ratio or Maximum Drawdown would improve the practical relevance and comparability of predictive models (Section 4.1).
+Crisis Periods Few studies include data from financial crises such as 2008–2009 or 2020–2021, focusing instead on stable periods.
+Including data from volatile periods can support stress-testing and help build models more resilient to real-world fluctuations (Section 4.2).
+Data Selection and Bias Most studies use English-language datasets and U.S.
+financial sources.
+While there is growing diversity in the types of data used (e.g., news, filings, social media), much of it is static, with limited updates or adaptation to changing financial contexts.
+Developing multilingual and globally representative datasets (Section 4.4), along with maintaining and updating existing ones (Section 4.3), would support better generalization and long-term applicability of models across financial domains.
+PLM/LLM Adoption General-purpose PLMs and LLMs have largely replaced finance-specific and custom architectures.
+At the same time, the use of statistical NLP and conventional ML continues to decline.
+Revisiting finance-specific models and exploring alternative methods such as graph-based learning or hybrid statistical models could offer improvements for financial tasks (Sections 5.1, 5.2).
+Open Accessibility Code and dataset sharing has become more common, especially after 2021, enhancing transparency and reproducibility.
+Open access should be a standard in NLP-for-finance research when legally feasible.
+Maintaining functional and up-to-date repositories is key for long-term reproducibility (Section 5.3).
+This shift is also mirrored in the decline of custom architectures.
+Prior to 2022, task-specific models were more common: TagOp for TAT-QA (Zhu et al., 2021), HyBrider for HybridQA (Chen et al., 2020b)); MDRM (Qin and Yang, 2019), HTML (Yang et al., 2020a), and HAN (Hu et al., 2021) in volatility prediction (Niu et al., 2023; Mathur et al., 2022a)).
+Today, these efforts are rarely pursued, which seems like an efficiency gain at first glance.
+However, as Section 5.1 discusses, real-world financial applications often require design trade-offs, such as transparency for compliance, or fast inference for deployment, and off-the-shelf LLMs do not accommodate them well.
+The move away from custom models narrows the solution space at a time when financial tasks still demand methodological flexibility.
+5.3 Enabling Reproducibility through Open Resources
 As the adoption of language models accelerates, the issue of open accessibility in NLP research, especially in finance, becomes more relevant.
@@
 Maintaining functional, up-to-date repositories remains essential to support meaningful benchmarking and model development.
-7 Potential in Financial Forecasting
-Figure 6: Distribution of evaluation metrics used in Category I papers. Most rely on ML-based metrics, while only a few financial metrics appear repeatedly.
-Category I papers reveal several areas where forecasting models could be improved.
-Most studies use general ML metrics, leaving out finance-specific measures that better capture practical utility.
-Crisis periods are rarely considered, limiting insights into how models behave under stress.
-Finally, there is a strong U.S.
-and English-language bias in data, with limited adoption of global financial datasets.
-Expanding these areas of exploration could strengthen model robustness and support broader real-world applicability.
-7.1 Potential in Financial Metrics
-As shown in Figure 6, Category I papers mostly rely on ML metrics like accuracy, F1, or MSE (Sawhney et al., 2020b; Wu, 2020; Yangjia et al., 2022).
-While useful, these do not fully reflect financial performance.
-Only three financial metrics appear more than once – Sharpe Ratio, Maximum Drawdown, and Cumulative Return (defined in Appendix G).
-Finance-specific metrics offer risk-adjusted insights that ML measures overlook.
-Sharpe Ratio, for example, evaluates return relative to risk (Zou et al., 2022).
-Incorporating such metrics would improve comparability across models and align research outcomes with practical needs (Zhang et al., 2024; Zou et al., 2022; Sawhney et al., 2021b).
-7.2 Crisis Years for Stress Testing
-Figure 7: Distribution of data years used in Category I papers, highlighting publication trends, recession periods, and key financial data milestones.
-Figure 7 shows that data usage begins to rise after 1993, when electronic filings became publicly available, with another increase in 2005 following the HTML filing mandate (U.S.
-Securities and Exchange Commission, ).
-However, usage dips around 2009 and 2020 – years associated with the global financial crisis and COVID-19 pandemic.
-These periods are essential for evaluating model robustness, yet they remain underused, likely due to data availability challenges prior to 2015 (Bloomberg L.P., 2015).
-Despite these constraints, crisis periods offer valuable testing grounds.
-Most models are trained on stable conditions, which may not reflect their behavior during volatility.
-Including data from crises enables stress testing (Investopedia Staff, 2025b) and helps build forecasting models that remain reliable under instability.
-7.3 Data Sources
-Commonly Used and Overlooked Data Sources
-Most forecasting models use common financial sources: stock prices, SEC filings, financial news, earnings calls (e.g., MDRM (Qin & Yang, 2019), StockNet (Xu & Cohen, 2018)).
-But many valuable datasets are either rarely used (e.g., Federal Reserve reports (Menzio et al., 2024)) or completely absent from NLP research.
-Public resources like FRED (Federal Reserve Bank of St.
-Louis, 2024) and Fama-French (Fama & French, 2024) offer rich macroeconomic indicators that remain unexplored.
-Incorporating these datasets could help capture macroeconomic trends and market dynamics, improving model reliability.
-Data Accessibility and Maintenance
-Access limitations are also a concern.
-Legacy datasets like MDRM are outdated, and APIs such as Yahoo Finance’s have been discontinued.
-Others, like CRSP, require paid access.
-This limits reproducibility and restricts access to reliable, up-to-date data.
-There’s a clear need for open, regularly updated datasets applicable for financial forecasting.
-8 Discussion & Conclusion
-This paper presents a systematic, quantitative analysis of NLP research applied to finance.
-From an initial pool of 374 papers across major NLP venues (Appendix E), we curated a final set of 221 papers after filtering.
-We categorized them by primary task (Section 3) and assessed each across 11 qualitative and quantitative dimensions (Table 5).
-Table 3 summarizes key trends and identifies opportunities where language modeling research can further contribute to finance-related problems.
-Criteria Trends and Observations Potential Opportunities and Recommendations
-NLP Tasks in Finance Sentiment analysis, information extraction, and question answering are the most frequently addressed tasks.
-Forecasting is mostly centered around stock and volatility prediction.
-Underexplored areas such as explainability, privacy-preserving methods, and tasks like bankruptcy or cryptocurrency prediction present valuable directions for future work (Section 3).
-PLM/LLM Adoption General-purpose PLMs and LLMs have largely replaced finance-specific and custom architectures.
-At the same time, the use of statistical NLP and conventional ML continues to decline (Section 4).
-Revisiting finance-specific models and exploring alternative methods such as graph-based learning or hybrid statistical models could offer improvements for financial tasks (Section 5).
-Evaluation Metrics Most studies rely on ML metrics such as accuracy and MSE, which do not reflect financial performance.
-Incorporating financial metrics like Sharpe Ratio or Maximum Drawdown would improve the practical relevance and comparability of predictive models (Section 7.1).
-Crisis Periods Few studies include data from financial crises such as 2008–2009 or 2020–2021, focusing instead on stable periods.
-Including data from volatile periods can support stress-testing and help build models more resilient to real-world fluctuations (Section 7.2).
-Data Selection and Bias Most studies use English-language datasets and U.S.-based financial sources.
-While there is growing diversity in the types of data used (e.g., news, filings, social media), much of it is static, with limited updates or adaptation to changing financial contexts.
-Developing multilingual and globally representative datasets, along with maintaining and updating existing ones, would support better generalization and long-term applicability of models across financial domains (Section 7.3).
-Open Accessibility Code and dataset sharing has become more common, especially after 2021, enhancing transparency and reproducibility.
-Open access should be a standard in NLP-for-finance research when legally feasible.
-Maintaining functional and up-to-date repositories is key for long-term reproducibility (Section 5).
-Table 3: Summary of trends and research directions in NLP research applied to finance, highlighting areas of focus, methodological shifts, and key opportunities for future exploration.
-Implications for NLP & LM Researchers
-NLP research applied to finance has largely centered on a few core tasks – sentiment analysis, information extraction, and question answering – while underexploring others like explainability and privacy-preserving methods (Section 3).
-The rapid adoption of PLMs and LLMs has made it easier to apply language models, but has also led to reduced methodological variety (Section 4).
-Graph-based models, statistical techniques, and other alternatives remain underused despite their potential (Section 5).
-While open-source practices have become more common, researchers should be more careful about maintaining functional and up-to-date repositories to ensure long-term reproducibility and meaningful benchmarking (Section 5).
-In addition, adopting financial-domain metrics can make evaluations more meaningful and aligned with practical goals (Section 7.1).
-Including crisis periods in datasets (Section 7.2) and improving data selection practices (Section 7.3) can help ensure robustness and consistency.
-More diverse datasets, especially beyond English and U.S.
-markets, are also needed to broaden applicability (Section 7.3).
+6 Conclusion & Discussion
+Drawing on Table 3, our survey shows that sentiment analysis, information extraction, and forecasting have driven NLP in finance, yet critical gaps remain: finance-specific evaluation metrics, stress-testing on crisis data, and truly global, multilingual datasets.
+These findings underscore both the field’s progress and the work ahead to build resilient, transparent, and inclusive financial NLP solutions.
+Below, we discuss their implications for researchers and practitioners.
+Implications for NLP and Language Modeling Researchers
+The application of LMs to financial domains presents challenges that differ from traditional NLP benchmarks.
+Our findings suggest that domain-specific requirements, such as interpretability, latency, and compliance, make classical or hybrid approaches highly relevant, even amid the LLM wave.
+The lack of financial evaluation metrics, underuse of crisis-period data, and limited attention to multilingual and global financial corpora signal areas where NLP researchers can contribute meaningfully.
+Moreover, the scarcity of domain-adapted models points to missed opportunities for more effective domain transfer.
+Future work should prioritize robustness, transparency, and real-world relevance over benchmark performance alone.
 Considerations for Financial Practitioners
-While NLP methods are increasingly used in financial contexts, several challenges remain.
-Many models are trained on static or incomplete datasets, leading to biases that may affect predictions (Section 7.3).
-Evaluations often omit real-world considerations such as market volatility and financial risk (Sections 7.2, 7.1).
-Incorporating financial benchmarks and testing models under stress conditions could improve their practical value.
-Continued progress in this direction can better align research outcomes with decision-making needs in financial settings.
-Ethics Statement
-This study does not assess the merit or quality of individual papers.
-We do not suggest that any category or method is inherently better than others.
-Our goal is to map the volume and distribution of research applying NLP methods to finance, and to offer a foundation for further study, without making judgments about the value of specific approaches.
-Acknowledgments
-We did not receive any specific funding for this work. We thank Arnav Hiray and Michael Galarnyk for their comments.
-Appendix A Authorship and Funding
-Based on the analysis we conducted, we also made a binary classification of the following categories:
-  • Authors affiliated with academia
-  • Authors affiliated with Industry
-  • Funding from industry
-  • Funding from government sources
-  • Funding from academic sources
-Figure 8: This figure displays the correlation between authors from industrial and academic backgrounds within various types of funding such as academic, industrial and governmental.
-As seen in the figure 8, we see that there is no major correlation between funding and authorship.
-Thus it does not imply that if an author with either industrial or academic is present in the list of authors, there will necessarily be funding from the industry/academia/government.
-Category Count
-Papers with Academic and Industrial authors 78
-Papers with only Academic authors 118
-Papers with only Industrial authors 25
-Papers with Industrial Funding 49
-Papers with Governmental Funding 61
-Papers with Academic Funding 35
-Table 4: Summary of the count of authors and funding within each category.
-Appendix B Various Dimensions surveyed
-Exploration dimension Description
-Primary task The main task addressed in the paper, such as volatility prediction (Wang et al., 2024).
-Sub-Tasks Additional tasks that support the primary objective, like using sentiment analysis to enhance stock price predictions (Jain & Agrawal, 2024).
-Methodology Techniques applied, ranging from traditional machine learning to deep learning and large language models. See Section 5.
-Code and data availability Whether the paper provides open-source code or data, and whether those artifacts are accessible (e.g., active links).
-Contribution What type of contribution the paper has such as a dataset, framework, model, evaluation, and so on.
-Comparison/baselines The baselines the paper compares their work to.
-Data source The types of data used, including financial reports, social media, news articles, or other financial documents.
-Metrics Evaluation criteria, including standard ML metrics (accuracy, precision) and finance-specific metrics (Sharpe ratio).
-Data years The time periods covered by the data, such as crisis years or more stable financial periods.
-Authorship Indicates what affiliation the authors of the paper have (academic, industrial).
-Funding source The origin of the research funding (academic, industrial, or governmental) if any.
-Table 5: This table outlines the dimensions examined in our analysis of NLP research focused on the financial domain.
-Dimensions that led to remarkable findings are discussed in the main paper, while an analysis of authorship and funding is provided in Appendix A.
-Appendix C Common Methods used
-Methodology Description
-Statistical NLP Methods like TF-IDF, Bag-of-Words, and n-grams focused on extracting statistical patterns from financial text.
-Embeddings Techniques such as Word2Vec, GloVe, and custom embeddings designed to map financial terms into vector spaces, improving downstream tasks.
-Conventional ML Algorithms like Logistic Regression, Support Vector Machines (SVM), and Decision Trees, often used for classification and risk prediction.
-Deep Learning (DL) Custom neural network architectures tailored for specific financial tasks such as stock price forecasting.
-PLMs Encoder-only models (e.g., BERT) and Encoder-Decoder models (e.g., T5) are used for tasks like document summarization and information extraction.
-LLMs Models like GPT-3.5 and LLaMA-2 focus on text generation and understanding complex financial language.
-Statistical Modeling Involves correlation analysis, Granger causality, etc., to understand relationships in financial data.
-Graphs Uses graph structures to model interactions in financial systems.
-Knowledge Graphs Integrating structured knowledge within financial tasks to enhance model performance and reliability.
-Table 6: This table categorizes the key methodologies into nine groups.“PLMs” stands for Pretrained Language Models, while “LLMs” stands for Large Language Models.
-Appendix D Task Description
-The following Table 7 and Table 8 describe the task description with examples.
-Primary Task Examples & Insights
-Finance-Related (Categories IV)
-Explainable AI (XAI) refers to methods that make machine learning outputs more interpretable and less of a "black box." These papers show how small input changes can drastically affect model explanations (Sinha et al., 2021) and highlight the instability of tools like LIME in sensitive applications (Burger et al., 2023).
-In finance, where decisions must be traceable and justifiable, clear and reliable explanation methods are essential for building trust in tasks like fraud detection and risk assessment.
-Privacy-preserving methods protect sensitive data while still allowing model training and analysis.
-Examples include homomorphic encryption for text similarity (Kim et al., 2022), federated learning for distributed training without raw data exchange (Zhao et al., 2024d), and domain adaptation that shares only model parameters (Xiao et al., 2024b).
-These methods are especially important in finance, where institutions must work with private data while staying compliant and avoiding breaches.
-Numerical reasoning involves working with numbers and structured data to solve problems or develop algorithms.
-This includes solving math word problems using declarative knowledge (Roy & Roth, 2018), learning numeracy through number embeddings (Duan et al., 2021), and pretraining models for verifying tabular claims (Akhtar et al., 2023).
-In finance, good numerical reasoning supports tasks like market prediction and risk evaluation by improving how models interpret and reason about numbers.
-Financial Applications (Category III)
-Sentiment analysis helps assess market mood and forecast financial movements using sources like news, social media, and reports.
-Studies analyze opinions in reports and news (Rodriguez Inserte et al., 2023), investor sentiment on social media (Guo et al., 2023a), and tone in financial texts (Choe et al., 2023).
-As markets become more influenced by public opinion, sentiment analysis plays a growing role in trading and analysis.
-Information extraction (IE) and relation extraction focus on identifying entities and linking them to events or other entities.
-Examples include extracting financial events from documents (Zheng et al., 2019), signals in reports (Huang et al., 2023), and patterns in social media (Conforti et al., 2022).
-These tools help connect companies, individuals, and financial instruments to real-world developments (Liou et al., 2021), supporting real-time analysis and prediction.
-Question Answering (QA) enables systems to find specific information in financial documents, reports, and databases.
-QA models support quick access to relevant facts (Mavi et al., 2023), and recent work focuses on calibrating large language models (LLMs) for finance-specific tasks (Zhao et al., 2024c; Theuma & Shareghi, 2024; Addlesee, 2024).
-This area emphasizes not only building QA tools but also improving their accuracy for use in high-stakes financial contexts.
-Table 7: Overview of key primary tasks in Categories IV (Finance-Related) above the mid line and Category III (Financial Applications) below.
-Each entry summarizes the task’s role and relevance in NLP research applied to finance, with representative examples and practical insights.
-Primary Task Examples & Insights
-Financial Resources (Category II)
-Dataset and Resource Construction involves building structured, labeled financial datasets for model development and evaluation.
-These papers focus on creating large-scale annotated resources (Chen et al., 2020a), including datasets like Tab-CQA (Liu et al., 2023) and ConvFinQA (Chen et al., 2022c) designed for reasoning over tables and multi-step numerical queries.
-Other works assemble corpora from Earnings calls (Pardawala et al., 2025), financial reports (Shah et al., 2022), company filings (Zmandar et al., 2022), news (Tang et al., 2023), and government documents (Shah et al., 2023), enabling downstream tasks like entity linking, numerical reasoning, or forecasting.
-Fraud detection leverages domain-specific data and models to identify deceptive financial behavior across platforms.
-Erben & Waldis (2024) identifies financial scams on Instagram using a fine-tuned BERT model deployed via browser extension and REST API.
-Another approach detects identity fraud through interactive dialogue, employing knowledge graphs and reinforcement learning to expose inconsistencies in claimed personal data (Wang et al., 2019).
-These systems highlight how NLP architectures can protect users from financial manipulation.
-Financial Forecast (Category I)
-Stock Price and Volatility Prediction aims to forecast stock movements or market instability using text data.
-These models use historical data (Sawhney et al., 2021a), news (Ahbali et al., 2022), and event-driven signals (Sawhney et al., 2020a; Wu, 2020) to predict stock trends.
-Volatility prediction further explores market sensitivity by analyzing unstructured text like press releases or financial articles (Qin & Yang, 2019), helping traders anticipate fluctuations.
-Risk Prediction involves identifying the likelihood and impact of adverse financial events, such as defaults or market disruptions.
-These models analyze unstructured data like earnings call transcripts (Sang & Bao, 2022), regulatory filings, and legal documents to detect early signals of financial risk (Li et al., 2023b).
-Applications include credit risk estimation, fraud detection, and systemic risk monitoring (Zhang et al., 2024).
-Table 8: Descriptions of major primary tasks in Category II (Financial Resources) above the line and Category I (Financial Forecast) below.
-Each entry summarizes the task’s focus and contribution to NLP research applied to finance, with representative examples.
-Appendix E Conference Proceedings
-Category Conferences and Workshops
-Financial Forecast (Category I) ACL, Australasian Language Technology Association Workshop, CCL, EACL, EcoNLP Workshop, EMNLP, EMNLP (Industry), Financial Technology and Natural Language Processing (FinNLP) Workshop, ICCL, LREC-COLING, NAACL, NAACL (Industry), SRW
-Financial Resources (Category II) ACL, ACL (Industry), CoNLL, EACL, EcoNLP Workshop, EMNLP, EMNLP-IJCNLP, e-Commerce and NLP Workshop, Financial Narrative Processing and MultiLing Financial Summarisation Workshop, Financial Technology and Natural Language Processing (FinNLP) Workshop, GWC, ICON, ICCL, LREC, LREC-COLING, NAACL, NAACL (Industry), NLP4PI Workshop, Pattern-based Approaches to NLP in the Age of Deep Learning Workshop, SIGHUM Workshop
-Financial Applications (Category III) ACL, ATALA, Bridging Human–Computer Interaction and Natural Language Processing Workshop, CCL, CoLM, Computational Approaches to Subjectivity, Sentiment and Social Media Analysis Workshop, DeeLIO Workshop, EACL, EcoNLP Workshop, EMNLP, EMNLP (Industry), EMNLP-IJCNLP, Financial Narrative Processing and MultiLing Financial Summarisation Workshop, Financial Technology and Natural Language Processing (FinNLP) Workshop, ICON, ICCL, INLG, LREC, LREC-COLING, NAACL, NAACL (Industry), Natural Legal Language Processing Workshop, News Media Content Analysis and Automated Report Generation Workshop, Pattern-based Approaches to NLP in the Age of Deep Learning Workshop, Safety4ConvAI Workshop, Structured Prediction for Natural Language Processing Workshop, TextGraphs
-Finance Related (Category IV) ACL, ACL-IJCNLP, BlackboxNLP Workshop, EMNLP, Financial Technology and Natural Language Processing (FinNLP) Workshop, LREC-COLING, SRW, TrustNLP Workshop
-Table 9: Conference and Workshop Categorization
-Appendix F Papers per year
-Year 2024 2023 2022 2021 2020 2019 2018 2017 2016 2015 2014 2012 2010 2008 2006 2002 2001 1976 1975
-Number of Papers 127 53 90 32 57 15 14 21 6 1 5 1 3 3 1 1 1 1 1
-Table 10: Number of papers satisfying our filtration criterion per year.
-Appendix G Glossary
-Metric Definition
-Sharpe Ratio Measures how much return an investment gives for each unit of risk, compared to a risk-free asset (Investopedia Staff, 2025a).
-Maximum Drawdown The biggest drop from a peak to a low point in a portfolio’s value before it recovers (Investopedia Staff, 2023b).
-Cumulative Return The total profit or loss from an investment over time (Investopedia Staff, 2023a).
-Table 11: Key Definitions of financial terms used in our paper.
+For finance professionals and institutions considering NLP solutions, this survey provides guidance on selecting appropriate tools.
+Off-the-shelf LLMs may appear attractive, but their performance in finance-specific contexts is often limited by lack of customization, interpretability, or latency guarantees.
+Classical and lexicon-based methods, when tailored to regulatory or operational constraints, can outperform black-box models in compliance or auditing settings.
+Practitioners should also evaluate model performance using finance-aligned metrics and test under historical stress scenarios.
+Finally, leveraging multilingual datasets and broadening market coverage can help mitigate geographic and systemic biases in automated decision-making.

### v3

@@
 Abstract
-Recent advances in language modeling have led to growing interest in applying Natural Language Processing (NLP) techniques to financial problems, enabling new approaches to analysis and decision-making.
+Recent advances in language modeling have led to a growing number of papers related to finance in top-tier Natural Language Processing (NLP) venues.
 To systematically examine this trend, we review 374 NLP research papers published between 2017 and 2024 across 38 conferences and workshops, with a focused analysis of 221 papers that directly address finance-related tasks.
-We evaluate these papers across 11 quantitative and qualitative dimensions, and our study identifies the following opportunities: (i) expanding the scope of forecasting tasks; (ii) enriching evaluation with financial metrics; (iii) leveraging multilingual and crisis-period datasets; and (iv) balancing PLMs with efficient or interpretable alternatives.
-We identify actionable directions for research and practice, supported by dataset and tool recommendations, with implications for both the academia and industry communities.
+We evaluate these papers across 11 quantitative and qualitative dimensions, and our study identifies the following opportunities for NLP researchers: (i) expanding the scope of forecasting tasks; (ii) enriching evaluation with financial metrics; (iii) leveraging multilingual and crisis-period datasets; and (iv) balancing PLMs with efficient or interpretable alternatives.
+We identify actionable directions supported by dataset and tool recommendations, with implications for both the academia and industry communities.
+Language Modeling for the Future of Finance: A Survey into Metrics, Tasks, and Data Opportunities
+Nikita Tatarinov $\spadesuit$ 🖂, Siddhant Sukhani $\spadesuit$ 🖂, Agam Shah 🖂, Sudheer Chava Georgia Institute of Technology, Atlanta, USA $\spadesuit$  Indicates equal contribution 🖂  Corresponding Authors: {ntatarinov3, ssukhani3, ashah482}@gatech.edu
 1 Introduction
-Language modeling has become a key tool in natural language processing (NLP) for analyzing unstructured text, including reports, news articles, and social media posts.
-These methods are increasingly applied to finance-related tasks such as sentiment analysis (Balakrishnan et al., 2022), information extraction (Huang et al., 2023), summarization (Khanna et al., 2022), stock prediction (Jain and Agrawal, 2024), and volatility forecasting (Niu et al., 2023).
-Given the increasing number of NLP papers addressing financial problems, there is a need to systematically examine how the field is contributing to finance-related applications.
-While prior surveys have examined this intersection (Table 1), many lack a structured approach (Chen et al., 2022a; Gao et al., 2021b; Xiao et al., 2024), focusing on broad NLP techniques (Jagdale and Deshmukh, 2025; Man et al., 2019; Liu, 2024), deep learning (Ozbayoglu et al., 2020) and Large Language Models (LLMs) (Nie et al., 2024; Li et al., 2023), or specific tasks such as sentiment analysis (Mishev et al., 2020).
-In this study, we focus specifically on NLP research applied to finance.
-Our scope includes papers published in NLP venues where methods are developed, benchmarked, or adapted for financial problems.
-We analyze 374 papers published from 2017 to 2024 across 38 NLP conferences and workshops.
-After further filtering (Section 2), we retain 221 papers that directly address finance-related tasks.
-These are evaluated across 11 quantitative and qualitative dimensions, including tasks, methodologies, datasets, evaluation metrics, and accessibility.
+Language modeling is a core method in natural language processing (NLP) for analyzing unstructured text Peters et al.
+(2018); Brown et al.
+(2020).
+At the same time, finance has become one of NLP’s primary application domains: as seen in Figure 5, the number of finance-related papers in top-tier NLP venues has been rising rapidly year over year.
+The tasks in these papers span general NLP problems on financial data, including sentiment analysis (Balakrishnan et al., 2022), information extraction (Huang et al., 2023), and summarization (Khanna et al., 2022), as well as finance problems addressed with NLP techniques, such as stock prediction (Jain and Agrawal, 2024), volatility forecasting (Niu et al., 2023), etc.
+The primary goal of our study is to highlight the opportunities to advance NLP research when applied to finance.
+Our scope includes 374 papers published from 2017 to 2024 across 38 NLP conferences and workshops.
+After further filtering (Section 2), we retain 221 papers that directly address finance-related tasks and evaluate them across 11 quantitative and qualitative dimensions, including tasks, methodologies, datasets, evaluation metrics, and accessibility.
+To the best of our knowledge, we present the first study to have a systematic look at finance-domain research in NLP venues, avoiding selection bias: existing surveys (Table 1) manually select the reviewed papers.
+In addition, some works are focusing on specific NLP approaches such as deep learning (Ozbayoglu et al., 2020) and Large Language Models (LLMs) (Nie et al., 2024; Li et al., 2023b), or on specific tasks such as sentiment analysis (Mishev et al., 2020).
 Table 1: Comparison of previous Natural Language Processing surveys in finance based on their focus areas, number of papers reviewed, and analytical features.
@@
 Paper Review Years Papers Reviewed Features Area surveyed Systematic Collection Domain trends Quantitative Analysis Temporal Analysis
-(Gao et al., 2021a) 1959 (A)-2020 87 5 General Overview × × × ×
+(Gao et al., 2021) 1959 (A)-2020 87 5 General Overview × × × ×
 (Liu, 2024) 2022-2024 49 11 General Overview × × × ×
 (Millo et al., 2024) 2018-2023 30 1 Methodologies × × ×
-(Chen et al., 2020a) 2016-2019 62 3 Financial Technology × × × ×
-(Chen et al., 2022b) 2018-2022 38 2 Financial Technology × × ×
+(Chen et al., 2020b) 2016-2019 62 3 Financial Technology × × × ×
+(Chen et al., 2022a) 2018-2022 38 2 Financial Technology × × ×
 (Xing et al., 2017) 1998 (A)-2016 127 4 Financial Forecasting × × ×
-(Zhao et al., 2024) 2004 (A)-2024 146 1 Large Language Models × × ×
+(Zhao et al., 2024a) 2004 (A)-2024 146 1 Large Language Models × × ×
 (Li et al., 2024) 2020-2023 68 1 Large Language Models × × × ×
@@
 Our analysis reveals not only valuable insights but also actionable directions for both research and practice with dataset and tool recommendations.
-Forecasting tasks, while well-established, leave room for exploration, especially in areas such as risk and macroeconomic prediction (Section 3).
+Financial forecasting tasks, while well-established, leave room for exploration, especially in areas such as risk and macroeconomic prediction (Section 3).
 Financial evaluation metrics are gaining traction and could further improve the practical applicability of models (Section 4.1).
 The growing availability of temporally diverse (Sections 4.2, 4.3), multilingual and multimodal (Sections 4.4) datasets enables more robust, globally applicable models.
-Finally, while PLMs and LLMs dominate, lightweight and interpretable methods remain viable for practical deployment (Section 5).
 2 Paper Extraction Process
 To study how NLP has been applied to finance, we selected papers from 38 NLP venues, including ACL, NAACL, EMNLP, LREC, CoLM, COLING, and workshops like FinNLP and the Workshop on Economics and Natural Language Processing.
-As illustrated in Figure 1, we began by filtering papers that mentioned “finance” or “financial” in their abstracts.
-Out of all such papers published from 1975 to the present, more than 94% were published from 2017 onward.
-This growth coincides with the emergence of transformer-based models (Vaswani et al., 2017), which greatly expanded the scope of NLP applications.
-Accordingly, we use 2017 as a threshold year for our analysis.
-The keyword filtering approach enabled us to cast a wide net, allowing for a high volume of papers.
-We were then able to filter these further using manual methods.
+As illustrated in Figure 1, we began by filtering papers that mentioned “finance” or “financial” in their abstracts Mackenzie et al.
+(2018); Nogueira and Lin (2020), enabling us to cast a wide net, allowing for a high volume of papers that could be manually postprocessed.
+Out of all such papers published from 1975 to the present, more than 94% appeared from 2017 onward, coinciding with the emergence of transformer-based models (Vaswani et al., 2017), which greatly expanded the scope of NLP applications Devlin et al.
+(2019); Dai et al.
+(2019); Lewis et al.
+(2020), and we therefore use 2017 as a threshold year for our analysis.
 This initial filtering returned 374 papers.
@@
 After removing these, we retained 221 papers for our analysis.
-Unlike previous surveys, which often emphasize qualitative observations, our study combines quantitative and qualitative methods for a broader view of the field.
-We categorized the selected papers by their primary focus into four task-based categories (Figure 2).
+Unlike previous surveys as shown in Table 1, which often emphasize qualitative observations Gao et al.
+(2021); Liu (2024); Millo et al.
+(2024), our study combines quantitative and qualitative methods for a broader view of the field.
+In comparison to other works, we categorized the selected papers by their primary focus into four task-based categories (Figure 2).
 This classification helps us examine trends in how NLP techniques are used to tackle finance-relevant problems and supports more detailed comparisons across types of tasks.
@@
 To understand how NLP has been used in financial contexts, we categorized papers into four distinct groups (shown in Table 2) based on their primary tasks (Figure 2).
-This classification provides a mapping from language modeling techniques to their applications and the level of engagement with finance, with Category I papers representing direct financial use cases.
 Table 2: Summary of task categories based on the specifics of their financial focus, ranging from direct prediction of financial outcomes to general NLP tasks with potential financial value.
@@
 In the Financial Applications (Category III) group, sentiment analysis (Rodriguez Inserte et al., 2023), information extraction (Lior et al., 2024), and question answering (Kosireddy et al., 2024) are the dominant tasks.
-These methods are widely used to analyze earnings calls, reports, and market commentary, where investor sentiment and factual extraction are key inputs for decision-making.
+These methods are widely used to analyze earnings calls, reports, and market commentary, where investor sentiment and factual extraction are key inputs for decision-makingChen et al.
+(2021); Zhu et al.
+(2021); Mukherjee et al.
+(2022); Qin and Yang (2019).
 Recent efforts have aimed to improve QA models for handling financial text (Theuma and Shareghi, 2024; Liu et al., 2024; Mavi et al., 2023).
@@
 Numerical reasoning (Akhtar et al., 2023), important for understanding financial statements and assessing risk, is receiving more attention.
-However, other areas – such as explainable AI (XAI) (Klein and Walther, 2024) and privacy-preserving methods (Abbe et al., 2012) – remain rarely explored, despite their relevance to secure and interpretable decision-making (Basu et al., 2021).
+However, other areas – such as explainable AI (XAI) (Klein and Walther, 2024) and privacy-preserving methods (Abbe et al., 2012) remain rarely explored, despite their relevance to secure and interpretable decision-making (Basu et al., 2021).
 Figure 2: Distribution of primary tasks across categories.
@@
 4.1 Advancing Evaluation with Finance Specific Metrics
-As shown in Figure 3, Category I papers mostly rely on ML metrics such as accuracy, F1, or MSE (Sawhney et al., 2020; Wu, 2020; Yangjia et al., 2022).
+As shown in Figure 3, Category I papers mostly rely on ML metrics such as accuracy, F1, or MSE (Sawhney et al., 2020b; Wu, 2020; Yangjia et al., 2022).
 While useful, these do not fully reflect financial performance.
 Finance-specific metrics offer risk-adjusted insights that ML measures overlook.
-Incorporating such metrics would improve comparability across models and align research outcomes with practical needs (Zhang et al., 2024; Zou et al., 2022; Sawhney et al., 2021).
-Consider two models that predict daily stock returns.
-Model A achieves a lower mean squared error (MSE) than Model B, yet when deployed in a trading strategy, it results in frequent trades, high volatility, and poor performance during market downturns.
-In contrast, Model B, despite its higher MSE, generates a smoother equity curve, experiences lower drawdown, and delivers a higher Sharpe Ratio thus making it more appealing from an application perspective.
-This highlights a critical insight emphasized in financial machine learning literature: traditional statistical metrics like MSE are often inadequate for evaluating trading strategies, as they fail to capture real-world profitability and risk exposure (Bailey et al., 2015; de Prado, 2018).
+Incorporating such metrics would improve comparability across models and align research outcomes with practical needs (Zhang et al., 2024; Zou et al., 2022; Sawhney et al., 2021c).
+As highlighted by Tamar et al.
+(2012); Liu et al.
+(2022), the financial machine learning literature stresses that conventional statistical metrics such as mean squared error are insufficient for assessing trading strategies, since they overlook the dimensions of real-world profitability and risk management (Bailey et al., 2015; de Prado, 2018).
 Instead, financial metrics such as Sharpe Ratio, Maximum Drawdown, and Cumulative Return are essential because they reflect the model’s risk-adjusted performance and robustness across market regimes (Lo, 2002; Krauss et al., 2017; Takahashi et al., 2009) as well as the applicability of the model to practical real world situations.
-Key Financial Metrics
-There are certain financial metrics that measure models’ performance in a financial setting and provide insight into the their applicability.
-Sharpe Ratio quantifies risk-adjusted return by comparing excess returns to volatility (Investopedia, 2025a).
-Maximum Drawdown (MDD) captures the largest loss from peak to trough, indicating downside risk and model robustness (Hayes, 2024).
-Cumulative Return measures total profit over a period: a direct benchmark for model outputs (Investopedia, 2007).
-Tools for Metric Calculation
-Several Python libraries support quantitative finance and model-driven investment strategies.
-PyPortfolioOpt (Martin, 2021) offers portfolio optimization tools including Sharpe Ratio and risk-return metrics.
-QuantLib (QuantLib Team, 2000) provides analytics for pricing, drawdown, and custom risk modeling.
-Backtrader (Backtrader Team, 2015) enables backtesting of trading strategies with performance evaluation, ideal for simulating NLP-driven trading.
-bt (Morissette, 2025) supports rapid prototyping and comparison of forecasting models with built-in financial metrics.
+Key Financial Metrics and Python Libraries
+In financial machine learning, specialized metrics are essential for evaluating a model’s real-world applicability.
+The Sharpe Ratio quantifies risk-adjusted return by comparing excess returns to volatility (Investopedia, 2025a), while Maximum Drawdown (MDD) captures the largest peak-to-trough loss, reflecting downside risk and robustness (Hayes, 2024).
+Similarly, Cumulative Return provides a direct benchmark by measuring total profit over a period (Investopedia, 2007).
+Several Python libraries facilitate the computation of these metrics within model-driven investment workflows: PyPortfolioOpt (Martin, 2021) offers portfolio optimization tools including Sharpe and risk-return analysis; QuantLib (QuantLib Team, 2000) supports pricing, drawdown, and advanced risk modeling; Backtrader (Backtrader Team, 2015) enables backtesting and performance evaluation of trading strategies, including NLP-driven pipelines; and bt (Morissette, 2025) streamlines prototyping and comparison of forecasting models with built-in financial metrics.
 4.2 Strengthening Robustness by Incorporating Crisis Periods
@@
 SEC, 2025).
-However, most publications use post-2009 (the global financial crisis) and pre-2020 (the COVID-19 pandemic) data, thus missing the crisis periods in evaluation.
-Crisis periods are essential for evaluating model robustness, allowing stress testing (Investopedia, 2025b) and helping build forecasting models that remain reliable under instability.
-Consider the case of Long-Term Capital Management (LTCM).
-Its quantitative models were highly leveraged and calibrated almost exclusively on stable market data, failing to account for extreme tail events.
-When the 1997 Asian and 1998 Russian crises struck, LTCM’s Value-at-Risk (VaR) models severely underestimated exposure, leading to large-scale losses as spreads exploded and liquidity evaporated.
-Within weeks, LTCM lost over 90% of its equity, triggering systemic risk concerns across the global financial system (Department of Land Economy, 2003; Federal Reserve History, 2025; Vanity Fair Staff, 1998).
-As noted in academic and regulatory analyses, this collapse illustrates the dangers of relying solely on backtests from stable periods without incorporating the possibility of rare but catastrophic shocks.
-Most current models are trained in stable conditions that may not reflect their behavior during volatility.
+However, most publications use post-2009 (the global financial crisis) and pre-2020 (the COVID-19 pandemic) data, thus missing the crisis periods in evaluation Kim et al.
+(2022b); Fan et al.
+(2023); Yao et al.
+(2022).
+Crisis periods are essential for evaluating model robustness, allowing stress testing (Investopedia, 2025b; Hafiz et al., 2023; Aldasoro et al., 2025) and helping build forecasting models that remain reliable under instability.
+For example within industrial applications of such models, LTCM’s overdependence on stable market data and highly leveraged VaR models left it blind to tail risks, so when the Asian and Russian crises hit, spreads exploded, liquidity vanished, and the fund lost over 90% of its equity, sparking systemic risk concerns (Department of Land Economy, 2003; Federal Reserve History, 2025; Vanity Fair Staff, 1998).
 While there were data availability challenges prior to 2015 (Bloomberg L.P., 2015), the crisis years remain underused.
@@
 Most forecasting models use common financial sources: stock prices, SEC filings, financial news, earnings calls.
-But many valuable datasets are either rarely used (e.g., Federal Reserve reports (Shah et al., 2023; Menzio et al., 2024)) or completely absent from NLP research (e.g., shareholder letters).
+But many valuable datasets are either rarely used (e.g., Federal Reserve reports (Shah et al., 2023; Menzio et al., 2024a)) or completely absent from NLP research (e.g., shareholder letters).
 Public resources like FRED (Federal Reserve Bank of St.
 Louis, 2024) and Fama-French (Fama and French, 2024) offer rich but underexplored macro indicators.
-Incorporating these datasets could help capture macroeconomic trends and market dynamics, improving model reliability.
-Below, we highlight sources of valuable financial data, including those already used by the community and the other beneficial ones.
+Incorporating these datasets into ablation studies and further analysis could help align the various findings with macroeconomic trends and market dynamics, improving model reliability Chakraborty et al.
+(2016); Xu and Cohen (2018a); Sawhney et al.
+(2020b).
+We highlight sources of valuable financial data, including those already used by the community and the other beneficial ones in Appendix G.
 Figure 3: Distribution of evaluation metrics used in Category I papers. Most rely on ML-based metrics, while only a few financial metrics appear repeatedly.
-Regulatory Filings and Earnings Calls
-U.S.
-public companies file a range of documents such as 10-K, 10-Q, 8-K, credit agreements, proxy statements, S-1, S-3, and others are accessible via the SEC’s EDGAR platform, through free API calls and RSS feeds (U.S.
-SEC, 2024).
-While similar registries exist (e.g., Canada’s SEDAR, Japan’s EDINET, UK’s Companies House), they are rarely used, reinforcing a U.S.-centric and English-language bias (Section 4.4).
-Earnings calls, where firms discuss quarterly performance with investors, are another rich data source.
-Seeking Alpha transcribes  4,500 calls quarterly (Seeking Alpha, 2025); other providers include The Motley Fool (The Motley Fool, 2025), FactSet (FactSet Insight, 2025), Thomson Reuters (Thomson Reuters, 2025), and S&P Global (S&P Global Market Intelligence, 2025), though often behind paywalls.
-Companies also post transcripts/slides on IR websites.
-Public datasets include transcriptions and audio, such as S&P 500 calls from 2017 (Qin and Yang, 2019) and 1,213 companies’ data covering 2015-2018 (Li et al., 2020a).
+4.4 Data Concerns and Biases
 Figure 4: Data year distribution in financial forecasting papers, annotated with major financial events and infrastructure milestones.
 Highlights underuse of crisis periods despite their importance for model robustness.
-Financial News and Social Media
-Real-time news from sources like Reuters, Bloomberg, Dow Jones, CNBC, FT, and WSJ is vital for forecasting.
-The Reuters-21578 dataset (UC Irvine, 2023), with 10,369 articles from 1987, remains a benchmark in text analysis.
-Larger archives such as FNSPID (Zhou et al., 2024) link 15.7 million articles (1999-2023) to S&P 500 stocks.
-News is accessible via APIs (e.g., Bloomberg (Bloomberg L.P., 2025a)) and free sources like Yahoo Finance (Yahoo Finance, 2025) and Google News (Google News, 2025).
-On social media, “FinTwit” and platforms like X and StockTwits are used for sentiment analysis.
-Key datasets include Sentiment140 (Kaggle Contributor, 2025; Chaudhary, 2020), StockNet (Xu and Cohen, 2018), and Reddit datasets (Kaggle Contributor, 2021; Wang et al., 2024).
-X’s API offers limited free and full historical access under paid tiers (X (2025), formerly Twitter).
-StockTwits, with Bullish/Bearish sentiment tagging, offers free and commercial API access (StockTwits, 2025) and is often used in forecasting studies.
-Macroeconomic and Monetary Data
-Central bank communications (policy statements, minutes, and press releases) offer critical macroeconomic guidance.
-The U.S.
-Federal Reserve’s FOMC releases detailed statements and transcripts on interest rates, inflation, and economic outlook (Federal Reserve System, 2025), with archives available on the Fed’s website.
-Despite their market relevance, such texts remain underused in NLP-based forecasting.
-Shah et al.
-(2023) addresses this by introducing an annotated dataset of FOMC communications and proposing a sentence-level classification task (hawkish, dovish, neutral) using fine-tuned transformers.
-Economic indicators like GDP, CPI, and unemployment figures – issued by the BEA (U.S.
-BEA, 2025), FRED (Federal Reserve Bank of St.
-Louis, 2024), Eurostat, IMF, OECD, and the World Bank (The World Bank, 2025) – are mostly numerical but often include press commentary, all freely available via official portals.
-Analyst Reports
-These are proprietary equity research reports with buy/sell calls and sector outlooks.
-Data providers such as Thomson Reuters, FactSet, Bloomberg, and S&P I/B/E/S aggregate these reports, typically through subscription services.
-All-in-one platforms like the Bloomberg Terminal (Bloomberg L.P., 2025b) and S&P Capital IQ (S&P Global, 2025) provide access to equity research, earnings estimates, and corporate event data as part of their premium offerings.
-There is no comprehensive free source; however, excerpts or key takeaways occasionally surface in news articles or on platforms like Seeking Alpha.
-4.4 Data Concerns and Biases
 Data Accessibility Concerns
@@
 Other essential databases, such as CRSP, require paid access and are unavailable to most individual researchers.
-These barriers severely limit reproducibility and restrict access to reliable, up-to-date data.
-As a result, there is a clear need for open, regularly updated financial datasets suitable for forecasting.
+These barriers severely limit reproducibility and access to reliable data, underscoring the need for open, regularly updated financial datasets for forecasting.
 Data Language and Country Bias
 As shown in the previous sections, financial NLP research is heavily skewed toward English-language data, particularly from U.S.
-markets.
+markets (Chen et al., 2021; Reddy et al., 2024; Mukherjee et al., 2022).
 Most forecasting datasets rely on U.S.
@@
 contexts.
-Figure 6 highlights that while some multilingual datasets exist, they are mostly designed for language modeling rather than financial forecasting.
+Figure 7 highlights that while some multilingual datasets exist, they are mostly designed for language modeling rather than financial forecasting Ding et al.
+(2014); Cortis et al.
+(2017).
 Very few datasets enable forecasting tasks in other languages, with rare exceptions like the bilingual CMIN corpus (Luo et al., 2023).
@@
 Multimodal Data Scarcity
-Despite significant progress in multimodal learning, the domain of NLP in Finance still lacks comprehensive datasets that integrate text, audio, and other critical financial modalities.
-One of the few existing resources is MAEC, a dataset aligning transcripts with audio from over 3,400 S&P 500 earnings calls, totaling more than 920 hours of speech, specifically designed for financial risk forecasting (Li et al., 2020b).
-More recently, FinAudio has emerged as the first standardized benchmark focusing on audio-first tasks (ASR on short and long earnings call segments, plus summarization), compiling over 400 hours of financial audio from multiple sources to evaluate Audio-LLM capabilities (Cao et al., 2025).
-While many existing datasets support exploration of acoustic and textual cues in financial forecasting, truly multimodal resources (those integrating visual data such as stock charts, presentation slides alongside structured financial metrics) remain very limited.
+Despite significant progress in multimodal learning, the domain of NLP in Finance still lacks comprehensive datasets that integrate text, audio, and other critical financial modalities Sawhney et al.
+(2020b, 2021a); Kaikaus et al.
+(2022).
+MAEC, a dataset aligning transcripts with audio from over 3,400 S&P 500 earnings calls (920+ hours of speech) for financial risk forecasting, and FinAudio, the first standardized benchmark for audio-first tasks with 400+ hours of financial audio for evaluating Audio-LLM capabilities, represent two of the key recent resources in financial audio research (Li et al., 2020a; Cao et al., 2025).
+Our survey indicates that while many existing datasets support exploration of acoustic and textual cues in financial forecasting, truly multimodal resources (those integrating visual data such as stock charts, presentation slides alongside structured financial metrics) remain very limited.
 Notable exceptions include Mathur et al.
@@
 (2024), which begin to address this gap.
-This modality gap highlights the pressing need for richer, tri- or quad-modal datasets to better capture the multifaceted signals that influence market dynamics.
+This modality gap highlights the pressing need for richer, tri- or quad-modal datasets to better capture the multifaceted signals that influence market dynamics Liang and et al.
+(2022); Li and et al.
+(2024).
 5 Broader Insights for NLP in Finance
 5.1 Complementing Foundation Models with Practical Alternatives
-Based on our observations from our corpus of research papers, the methodological landscape in NLP for Finance has shifted decisively toward pretrained language models (PLMs) and large language models (LLMs).
+Based on our observations from our vast corpus of research, the methodological landscape in NLP for Finance has shifted decisively toward pretrained language models (PLMs) and large language models (LLMs).
 We group existing approaches into several broad categories, which allows us to track adoption trends and the rise of PLMs, beginning with the use of BERT-based models in finance around 2019 (Araci, 2019a; Yang et al., 2020c), and accelerating rapidly with the emergence of general-purpose LLMs like ChatGPT (Wu et al., 2023; Gao et al., 2023).
-However, NLP tasks in Finance often pose domain-specific constraints such as interpretability, regulatory transparency, data scarcity, and the need for low-latency systems.
+However, NLP tasks in Finance often pose domain-specific constraints such as interpretability, regulatory transparency, data scarcity, and the need for low-latency systems Luo et al.
+(2018); Maia et al.
+(2018).
 These factors make non-PLM methods (such as rule-based classifiers, feature-engineered models, and domain-specific embeddings) not only viable but often preferable in production environments (Shah et al., 2023; Lopez et al., 2021).
@@
 Latency and Scalability
-Classical NLP pipelines are far more resource-efficient than LLMs.
+Classical NLP pipelines are far more resource-efficient than LLMs Strubell et al.
+(2019).
 In time-time contexts like news analytics or alert systems, lightweight models (e.g., logistic regression or small CNNs) deliver insights instantly and can scale to massive corpora without requiring GPU infrastructure (Zhai et al., 2019).
@@
 Interpretability and Compliance
-Lexicon- and rule-based methods remain essential in regulated settings.
+Lexicon- and rule-based methods remain essential in regulated settings Moreno-Ortiz et al.
+(2020); Du et al.
+(2023).
 Financial sentiment lexicons (e.g., Loughran-McDonald) (Loughran and McDonald, 2011) and keyword rules are transparent, auditable, and often outperform neural models in compliance or risk-flagging tasks (Hosseini et al., 2018).
@@
 Information Extraction and Knowledge Graphs
-Named-entity and relation extraction systems power knowledge graphs and schema-driven applications.
+Named-entity and relation extraction systems power knowledge graphs and schema-driven applications Elhammadi et al.
+(2020); Hamad et al.
+(2024).
 These systems offer precision, structure, and explainability, which are ideal for fraud detection, Know Your Customer (KYC), and regulatory alignment, where PLMs often lack controllability (Szarvas et al., 2007; Tarnopolski et al., 2019).
@@
 Data Efficiency and Domain Adaptability
-Feature-based pipelines using TF-IDF, static embeddings, and classical classifiers (SVMs, XGBoost) work well with limited data.
+Feature-based pipelines using TF-IDF, static embeddings, and classical classifiers (SVMs, XGBoost) work well with limited dataWang and Manning (2012).
 They enable explicit domain knowledge integration (such as tagging financial terms or numerical cues) for credit scoring, bankruptcy prediction, and event detection (Li et al., 2018; Khandani et al., 2021; Alanis et al., 2022).
-Financial firms often use TF-IDF (Simha, 2021) and XGBoost (Kamau, 2024) to build effective models with limited labeled data without requiring costly infrastructure.
+Financial firms often use TF‑IDF (Simha, 2021) and XGBoost (Kamau, 2024) to build effective models with limited labeled data without requiring costly infrastructure.
 Structure and Summarization
-Unsupervised techniques like LDA and extractive summarization remain valuable for discovering latent topics or summarizing dense financial documents.
+Unsupervised techniques like LDA and extractive summarization remain valuable for discovering latent topics or summarizing dense financial documents Filippova et al.
+(2009); Agrawal et al.
+(2021).
 These methods are fast, interpretable, and competitive on factual corpora like earnings calls or filings (Blei et al., 2003; Nenkova and McKeown, 2011).
-Companies like BlackRock, Bank of America (Phoenix Strategy Group, 2025), and American Express (Arora and Radhakrishnan, 2020) apply topic modeling and extractive summarization to earnings calls and filings, using unsupervised or minimally supervised methods for reliable and traceable outputs.
+Companies like Bank of America (Phoenix Strategy Group, 2025) and American Express (Arora and Radhakrishnan, 2020) apply topic modeling and extractive summarization to earnings calls and filings, using unsupervised or minimally supervised methods for reliable and traceable outputs.
 5.2 Implications of Shift Toward General-Purpose Models
@@
 As Figure 7 shows, even earlier versions of FinBERT (Araci, 2019b; Yang et al., 2020b; Liu et al., 2020; Huang et al., 2022) are used more than newer ones, and recent finance-tuned models like FinGPT (Yang et al., 2023) remain rare.
-This trend reflects how easy it has become to apply strong general models out of the box.
 For practitioners, this means many models are optimized for benchmarks, not for the realities of finance, where interpretability, latency, or regulation often matter more than raw accuracy.
 Figure 5: Trends in code and dataset availability, highlighting the shift toward open-source practices and the growing accessibility of NLP resources for finance.
-Figure 6: Timeline of PLM/LLM adoption in NLP research applied to finance, alongside key datasets by task type: C (custom models), QA (question answering), IE (information extraction), OL (other language), and FD (financial documents).
-For each model, the top date indicates release; the bottom, first usage in the surveyed papers.
 Table 3: Summary of trends and research directions in NLP research applied to finance, highlighting areas of focus, methodological shifts, and key opportunities for future exploration.
@@
 This shift is also mirrored in the decline of custom architectures.
-Prior to 2022, task-specific models were more common: TagOp for TAT-QA (Zhu et al., 2021), HyBrider for HybridQA (Chen et al., 2020b)); MDRM (Qin and Yang, 2019), HTML (Yang et al., 2020a), and HAN (Hu et al., 2021) in volatility prediction (Niu et al., 2023; Mathur et al., 2022a)).
-Today, these efforts are rarely pursued, which seems like an efficiency gain at first glance.
+Prior to 2022, task-specific models were more common: TagOp for TAT-QA (Zhu et al., 2021), HyBrider for HybridQA (Chen et al., 2020c)); MDRM (Qin and Yang, 2019), HTML (Yang et al., 2020a), and HAN (Hu et al., 2021) in volatility prediction (Niu et al., 2023; Mathur et al., 2022a)).
+Today, these efforts are rarely pursued.
 However, as Section 5.1 discusses, real-world financial applications often require design trade-offs, such as transparency for compliance, or fast inference for deployment, and off-the-shelf LLMs do not accommodate them well.
@@
 5.3 Enabling Reproducibility through Open Resources
-As the adoption of language models accelerates, the issue of open accessibility in NLP research, especially in finance, becomes more relevant.
+As we witness an accelerated adoption of language models, the issue of open accessibility in NLP research, especially in finance, becomes more relevant.
 As shown in Figure 5, open-source practices have become more common in recent years.
@@
 6 Conclusion & Discussion
-Drawing on Table 3, our survey shows that sentiment analysis, information extraction, and forecasting have driven NLP in finance, yet critical gaps remain: finance-specific evaluation metrics, stress-testing on crisis data, and truly global, multilingual datasets.
+Drawing on Table 3, our survey shows that sentiment analysis, information extraction, and forecasting have driven NLP in finance, yet critical gaps remain: finance‑specific evaluation metrics, stress‑testing on crisis data, and truly global, multilingual datasets.
 These findings underscore both the field’s progress and the work ahead to build resilient, transparent, and inclusive financial NLP solutions.
-Below, we discuss their implications for researchers and practitioners.
 Implications for NLP and Language Modeling Researchers
 The application of LMs to financial domains presents challenges that differ from traditional NLP benchmarks.
-Our findings suggest that domain-specific requirements, such as interpretability, latency, and compliance, make classical or hybrid approaches highly relevant, even amid the LLM wave.
+Our findings suggest that domain-specific requirements, such as interpretability, latency, and compliance, make classical or hybrid approaches highly relevant.
 The lack of financial evaluation metrics, underuse of crisis-period data, and limited attention to multilingual and global financial corpora signal areas where NLP researchers can contribute meaningfully.
 Moreover, the scarcity of domain-adapted models points to missed opportunities for more effective domain transfer.
-Future work should prioritize robustness, transparency, and real-world relevance over benchmark performance alone.
+As a result, our observations indicate that future work should prioritize robustness, transparency, and real-world relevance over benchmark performance alone.
 Considerations for Financial Practitioners
@@
 Finally, leveraging multilingual datasets and broadening market coverage can help mitigate geographic and systemic biases in automated decision-making.
+Limitations
+For readability, we standardize task names and merge closely related variants when counting.
+When a paper spans multiple areas, we assign a single primary tag to keep statistics interpretable; a multi-tag alternative would yield slightly different totals.
+These are intentional choices to aid comparability.
+Ethical Considerations
+This study does not assess the merit or quality of individual papers.
+We do not suggest that any category or method is inherently better than others.
+Our goal is to map the volume and distribution of research applying NLP methods to finance, and to offer a foundation for further study, without making judgments about the value of specific approaches.
+Appendix A Timeline Figure
+This appendix includes the timeline Figure 7.
+Appendix B Authorship and Funding
+Based on the analysis we conducted, we also made a binary classification of the following categories:
+  • Academic authors
+  • Industrial authors
+  • Funding (industrial)
+  • Funding (governmental)
+  • Funding (academic)
+As seen in the Figure 6, we see that there is no major correlation between funding and authorship.
+Thus it does not imply that if an industrial/academic author is present in the list of authors, there will necessarily be funding from the industry/academia/government.
+This finding is surprising actually as one would expect that a paper written by only academics would have academic/governmental funding and the same for industrial authors and industrial funding.
+Table 4: Summary of Authors and Funding Categories
+Category Sum
+Academic and Industrial 78
+Only Academic authors 118
+Only Industrial authors 25
+Funding (industrial) 49
+Funding (governmental) 61
+Funding (academic) 35
+Appendix C Survey Dimensions, Methodologies, and Task Categories in Financial NLP Research
+Figure 6: This figure displays the correlation between authors from industrial and academic backgrounds within various types of funding such as academic, industrial and governmental.
+This appendix consolidates the key reference tables from our study, providing an overview of the dimensions examined (Table 5), the methodologies employed (Table 6), and the task categories identified in NLP research applied to the financial domain (Tables 7 and 8).
+Figure 7: Timeline of PLM/LLM adoption in NLP research applied to finance, alongside key datasets by task type: C (custom models), QA (question answering), IE (information extraction), OL (other language), and FD (financial documents).
+For each model, the top date indicates release; the bottom, first usage in the surveyed papers.
+Table 5: This table outlines the dimensions examined in our analysis of NLP research focused on the financial domain.
+Dimensions that led to significant findings are discussed in the main paper, while an analysis of authorship and funding is provided in Appendix B.
+Exploration dimension Description
+Primary task The main task addressed in the paper, such as volatility prediction (Wang et al., 2024a).
+Sub-Tasks Additional tasks that support the primary objective, like using sentiment analysis to enhance stock price predictions (Jain and Agrawal, 2024).
+Methodology Techniques applied, ranging from traditional machine learning to deep learning and large language models. See Section LABEL:sec:methodologies.
+Code and data availability Whether the paper provides open-source code or data, and the quality of this accessibility (e.g., active links).
+Contribution What type of contribution does the paper have (Dataset, framework, model, evaluation, etc.)?
+Comparability of research What other research did the researchers compare their work to?
+Data source The types of data used, including financial reports, social media, news articles, or other financial documents.
+Metrics Evaluation criteria, including standard ML metrics (accuracy, precision) and finance-specific metrics (Sharpe ratio).
+Data years The time periods covered by the data, such as crisis years or more stable financial periods.
+Authorship Indicates if the paper was authored or co-authored by industry professionals, reflecting practical applications.
+Funding source The origin of the research funding (academic, industrial, or governmental) and its influence on the research focus.
+Table 6: This table categorizes the key methodologies into nine groups.“PLMs” stands for Pretrained Language Models, while “LLMs” stands for Large Language Models.
+Methodology Description
+Statistical NLP Methods like TF-IDF, Bag-of-Words, and n-grams focused on extracting statistical patterns from financial text.
+Embeddings Techniques such as Word2Vec, GloVe, and custom embeddings designed to map financial terms into vector spaces, improving downstream tasks.
+Conventional ML Algorithms like Logistic Regression, Support Vector Machines (SVM), and Decision Trees, often used for classification and risk prediction.
+Deep Learning (DL) Custom neural network architectures tailored for specific financial tasks such as stock price forecasting.
+PLMs Encoder-only models (e.g., BERT) and Encoder-Decoder models (e.g., T5) are used for tasks like document summarization and information extraction.
+LLMs Models like GPT-3.5 and LLaMA-2 focus on text generation and understanding complex financial language.
+Statistical Modeling Involves correlation analysis, Granger causality, etc., to understand relationships in financial data.
+Graphs Uses graph structures to model interactions in financial systems.
+Knowledge Graphs Integrating structured knowledge within financial tasks to enhance model performance and reliability.
+Primary Task Examples & Insights
+Finance-Related (Categories IV)
+Explainable AI (XAI) refers to methods that make machine learning outputs more interpretable and less of a "black box." These papers show how small input changes can drastically affect model explanations (Sinha et al., 2021) and highlight the instability of tools like LIME in sensitive applications (Burger et al., 2023).
+In finance, where decisions must be traceable and justifiable, clear and reliable explanation methods are essential for building trust in tasks like fraud detection and risk assessment.
+Privacy-preserving methods protect sensitive data while still allowing model training and analysis.
+Examples include homomorphic encryption for text similarity (Kim et al., 2022a), federated learning for distributed training without raw data exchange (Zhao et al., 2024c), and domain adaptation that shares only model parameters (Xiao et al., 2024).
+These methods are especially important in finance, where institutions must work with private data while staying compliant and avoiding breaches.
+Numerical reasoning involves working with numbers and structured data to solve problems or develop algorithms.
+This includes solving math word problems using declarative knowledge (Roy and Roth, 2018), learning numeracy through number embeddings (Duan et al., 2021), and pretraining models for verifying tabular claims (Akhtar et al., 2023).
+In finance, good numerical reasoning supports tasks like market prediction and risk evaluation by improving how models interpret and reason about numbers.
+Financial Applications (Category III)
+Sentiment analysis helps assess market mood and forecast financial movements using sources like news, social media, and reports.
+Studies analyze opinions in reports and news (Rodriguez Inserte et al., 2023), investor sentiment on social media (Guo et al., 2023), and tone in financial texts (Choe et al., 2023).
+As markets become more influenced by public opinion, sentiment analysis plays a growing role in trading and analysis.
+Information extraction (IE) and relation extraction focus on identifying entities and linking them to events or other entities.
+Examples include extracting financial events from documents (Zheng et al., 2019), signals in reports (Huang et al., 2023), and patterns in social media (Conforti et al., 2022).
+These tools help connect companies, individuals, and financial instruments to real-world developments (Liou et al., 2021), supporting real-time analysis and prediction.
+Question Answering (QA) enables systems to find specific information in financial documents, reports, and databases.
+QA models support quick access to relevant facts (Mavi et al., 2023), and recent work focuses on calibrating large language models (LLMs) for finance-specific tasks (Zhao et al., 2024b; Theuma and Shareghi, 2024; Addlesee, 2024).
+This area emphasizes not only building QA tools but also improving their accuracy for use in high-stakes financial contexts.
+Table 7: Overview of key primary tasks in Categories IV (Finance-Related) above the mid line and Category III (Financial Applications) below.
+Each entry summarizes the task’s role and relevance in NLP research applied to finance, with representative examples and practical insights.
+Primary Task Examples & Insights
+Financial Resources (Category II)
+Dataset and Resource Construction involves building structured, labeled financial datasets for model development and evaluation.
+These papers focus on creating large-scale annotated resources (Chen et al., 2020a), including datasets like Tab-CQA (Liu et al., 2023) and ConvFinQA (Chen et al., 2022b) designed for reasoning over tables and multi-step numerical queries.
+Other works assemble corpora from financial reports (Shah et al., 2022), company filings (Zmandar et al., 2022), news (Tang et al., 2023), and government documents (Shah et al., 2023), enabling downstream tasks like entity linking, numerical reasoning, or forecasting.
+Fraud detection leverages domain-specific data and models to identify deceptive financial behavior across platforms.
+Erben and Waldis (2024) identifies financial scams on Instagram using a fine-tuned BERT model deployed via browser extension and REST API.
+Another approach detects identity fraud through interactive dialogue, employing knowledge graphs and reinforcement learning to expose inconsistencies in claimed personal data (Wang et al., 2019).
+These systems highlight how NLP architectures can protect users from financial manipulation.
+Financial Forecast (Category I)
+Stock Price and Volatility Prediction aims to forecast stock movements or market instability using text data.
+These models use historical data (Sawhney et al., 2021b), news (Ahbali et al., 2022), and event-driven signals (Sawhney et al., 2020a; Wu, 2020) to predict stock trends.
+Volatility prediction further explores market sensitivity by analyzing unstructured text like press releases or financial articles (Qin and Yang, 2019), helping traders anticipate fluctuations.
+Risk Prediction involves identifying the likelihood and impact of adverse financial events, such as defaults or market disruptions.
+These models analyze unstructured data like earnings call transcripts (Sang and Bao, 2022), regulatory filings, and legal documents to detect early signals of financial risk (Li et al., 2023a).
+Applications include credit risk estimation, fraud detection, and systemic risk monitoring (Zhang et al., 2024).
+Table 8: Descriptions of major primary tasks in Category II (Financial Resources) above the line and Category I (Financial Forecast) below.
+Each entry summarizes the task’s focus and contribution to NLP research applied to finance, with representative examples.
+Appendix D Conference Proceedings
+Table 9 categorizes key conferences and workshops according to the four primary task categories identified in our survey.
+Appendix E Papers per year
+Table 10 reports the number of papers per year from 1975 to 2024.
+Appendix F Glossary
+Table 11 provides definitions to the key financial metrics mentioned in this paper.
+Category Conferences and Workshops
+Financial Forecast (Category I) ACL, Australasian Language Technology Association Workshop, CCL, EACL, EcoNLP Workshop, EMNLP, EMNLP (Industry), Financial Technology and Natural Language Processing (FinNLP) Workshop, ICCL, LREC-COLING, NAACL, NAACL (Industry), SRW
+Financial Resources (Category II) ACL, ACL (Industry), CoNLL, EACL, EcoNLP Workshop, EMNLP, EMNLP-IJCNLP, e-Commerce and NLP Workshop, Financial Narrative Processing and MultiLing Financial Summarisation Workshop, Financial Technology and Natural Language Processing (FinNLP) Workshop, GWC, ICON, ICCL, LREC, LREC-COLING, NAACL, NAACL (Industry), NLP4PI Workshop, Pattern-based Approaches to NLP in the Age of Deep Learning Workshop, SIGHUM Workshop
+Financial Applications (Category III) ACL, ATALA, Bridging Human–Computer Interaction and Natural Language Processing Workshop, CCL, CoLM, Computational Approaches to Subjectivity, Sentiment and Social Media Analysis Workshop, DeeLIO Workshop, EACL, EcoNLP Workshop, EMNLP, EMNLP (Industry), EMNLP-IJCNLP, Financial Narrative Processing and MultiLing Financial Summarisation Workshop, Financial Technology and Natural Language Processing (FinNLP) Workshop, ICON, ICCL, INLG, LREC, LREC-COLING, NAACL, NAACL (Industry), Natural Legal Language Processing Workshop, News Media Content Analysis and Automated Report Generation Workshop, Pattern-based Approaches to NLP in the Age of Deep Learning Workshop, Safety4ConvAI Workshop, Structured Prediction for Natural Language Processing Workshop, TextGraphs
+Finance Related (Category IV) ACL, ACL-IJCNLP, BlackboxNLP Workshop, EMNLP, Financial Technology and Natural Language Processing (FinNLP) Workshop, LREC-COLING, SRW, TrustNLP Workshop
+Table 9: Conference and Workshop Categorization
+Year 2024 2023 2022 2021 2020 2019 2018 2017 2016 2015 2014 2012 2010 2008 2006 2002 2001 1976 1975
+Number of Papers 127 53 90 32 57 15 14 21 6 1 5 1 3 3 1 1 1 1 1
+Table 10: Number of papers satisfying our filtration criterion per year.
+Metric Definition
+Sharpe Ratio Measures how much return an investment gives for each unit of risk, compared to a risk-free asset (Investopedia, 2025a).
+Maximum Drawdown The biggest drop from a peak to a low point in a portfolio’s value before it recovers.
+Cumulative Return The total profit or loss from an investment over time.
+Table 11: Key Definitions of financial terms used in our paper.
+Appendix G Data Sources
+Document type Sources Accessibility (incl. access method) Modalities Update cadence
+Regulatory filings SEC EDGAR; other registries: SEDAR (Canada), EDINET (Japan), Companies House (UK) Public via web UI; free EDGAR APIs and RSS (U.S.
+SEC, 2024); other registries public but less commonly used Text/PDF Event-driven; periodic (annual/quarterly)
+Earnings calls (transcripts, slides, audio) Seeking Alpha (Seeking Alpha, 2025); The Motley Fool (The Motley Fool, 2025); FactSet (FactSet Insight, 2025); Thomson Reuters (Thomson Reuters, 2025); S&P Global (S&P Global Market Intelligence, 2025); company IR sites Mixed: many providers behind paywalls; IR sites often public; access via web UI and provider APIs (where offered) Text transcripts; audio Quarterly around earnings; event-driven (scheduled call dates)
+Financial news Reuters, Bloomberg, Dow Jones, CNBC, FT, WSJ; free aggregators like Yahoo Finance (Yahoo Finance, 2025) and Google News (Google News, 2025) Mixed: premium (e.g., Bloomberg) and free sources, including static datasets (UC Irvine, 2023; Zhou et al., 2024); access via web UIs and vendor APIs (e.g., Bloomberg API (Bloomberg L.P., 2025a)) Text (articles; headlines) Real-time / continuous
+Social media X (Twitter); Reddit; sentiment datasets for X (Kaggle Contributor, 2025; Chaudhary, 2020; Xu and Cohen, 2018b) and Reddit (Kaggle Contributor, 2021; Wang et al., 2024b) X API with limited free and paid tiers (X (2025), formerly Twitter); StockTwits free/commercial API (StockTwits, 2025); public web UIs Short-text posts; sentiment tags (StockTwits Bullish/Bearish) Real-time
+Macroeconomic & monetary communications U.S.
+Federal Reserve (FOMC) statements and transcripts (Federal Reserve System, 2025) Public via official web portals; archives available Text (policy statements, minutes, press releases) Scheduled (policy meetings, release calendars)
+Economic indicators BEA (U.S.
+BEA, 2025), FRED (Federal Reserve Bank of St.
+Louis, 2024), the World Bank (The World Bank, 2025) Public via official portals and APIs Numeric time series + accompanying text commentary Scheduled (monthly / quarterly)
+Analyst reports Bloomberg Terminal (Bloomberg L.P., 2025b); S&P Capital IQ (S&P Global, 2025) Paid/proprietary subscription platforms; enterprise access/APIs; occasional public excerpts in news/Seeking Alpha Text/PDF (reports, notes) Ongoing; event-driven (coverage initiations, earnings, sector updates)
+Table 12: Financial text/data sources by document type with sources, accessibility (and access method), modalities, and update cadence.
+Regulatory Filings and Earnings Calls
+U.S.
+public companies file a range of documents such as 10‑K, 10‑Q, 8‑K, credit agreements, proxy statements, S‑1, S‑3, and others are accessible via the SEC’s EDGAR platform, through free API calls and RSS feeds (U.S.
+SEC, 2024).
+While similar registries exist (e.g., Canada’s SEDAR, Japan’s EDINET, UK’s Companies House), they are rarely used, reinforcing a U.S.-centric and English-language bias (Section 4.4).
+Earnings calls, where firms discuss quarterly performance with investors, are another rich data source Mukherjee et al.
+(2022); Koval et al.
+(2023); Pardawala et al.
+(2025).
+Seeking Alpha transcribes  4,500 calls quarterly (Seeking Alpha, 2025); other providers include The Motley Fool (The Motley Fool, 2025), FactSet (FactSet Insight, 2025), Thomson Reuters (Thomson Reuters, 2025), and S&P Global (S&P Global Market Intelligence, 2025), though often behind paywalls.
+Companies also post transcripts/slides on IR websites.
+Public datasets include transcriptions and audio, such as S&P 500 calls from 2017 (Qin and Yang, 2019) and 1,213 companies’ data covering 2015-2018 (Li et al., 2020b).
+Financial News and Social Media
+Real-time news from sources like Reuters, Bloomberg, Dow Jones, CNBC, FT, and WSJ is vital for forecasting.
+The Reuters-21578 dataset (UC Irvine, 2023), with 10,369 articles from 1987, remains a benchmark in text analysis.
+Larger archives such as FNSPID (Zhou et al., 2024) link 15.7 million articles (1999-2023) to S&P 500 stocks.
+News is accessible via APIs (e.g., Bloomberg (Bloomberg L.P., 2025a)) and free sources like Yahoo Finance (Yahoo Finance, 2025) and Google News (Google News, 2025).
+On social media, “FinTwit” and platforms like X and StockTwits are used for sentiment analysis.
+Key datasets include Sentiment140 (Kaggle Contributor, 2025; Chaudhary, 2020), StockNet (Xu and Cohen, 2018b), and Reddit datasets (Kaggle Contributor, 2021; Wang et al., 2024b).
+X’s API offers limited free and full historical access under paid tiers (X (2025), formerly Twitter).
+StockTwits, with Bullish/Bearish sentiment tagging, offers free and commercial API access (StockTwits, 2025) and is often used in forecasting studies.
+Macroeconomic and Monetary Data
+Central bank communications (policy statements, minutes, and press releases) offer critical macroeconomic guidance Ahrens and McMahon (2021); Peskoff et al.
+(2023); Menzio et al.
+(2024b).
+The U.S.
+Federal Reserve’s FOMC releases detailed statements and transcripts on interest rates, inflation, and economic outlook (Federal Reserve System, 2025), with archives available on the Fed’s website.
+Despite their market relevance, such texts remain underused in NLP-based forecasting.
+Shah et al.
+(2023) addresses this by introducing an annotated dataset of FOMC communications and proposing a sentence-level classification task (hawkish, dovish, neutral) using fine-tuned transformers.
+Economic indicators like GDP, CPI, and unemployment figures – issued by the BEA (U.S.
+BEA, 2025), FRED (Federal Reserve Bank of St.
+Louis, 2024), Eurostat (Eurostat, 2025), IMF (International Monetary Fund, 2025), OECD (Organisation for Economic Co-operation and Development, 2025), and the World Bank (The World Bank, 2025) – are mostly numerical but often include press commentary, all freely available via official portals.
+Analyst Reports
+These are proprietary equity research reports with buy/sell calls and sector outlooks.
+Data providers such as Thomson Reuters, FactSet, Bloomberg, and S&P I/B/E/S aggregate these reports, typically through subscription services.
+All-in-one platforms like the Bloomberg Terminal (Bloomberg L.P., 2025b) and S&P Capital IQ (S&P Global, 2025) provide access to equity research, earnings estimates, and corporate event data as part of their premium offerings.
+There is no comprehensive free source; however, excerpts or key takeaways occasionally surface in news articles or on platforms like Seeking Alpha.